### v2

# Scaling Laws of Synthetic Data for Language Models

Abstract
Large language models (LLMs) achieve strong performance across diverse tasks, largely driven by high-quality web data used in pre-training. However, recent studies indicate this data source is rapidly depleting. Synthetic data emerges as a promising alternative, but it remains unclear whether synthetic datasets exhibit predictable scalability comparable to raw pre-training data. In this work, we systematically investigate the scaling laws of synthetic data by introducing SynthLLM, a scalable framework that transforms pre-training corpora into diverse, high-quality synthetic datasets. Our approach achieves this by automatically extracting and recombining high-level concepts across multiple documents using a graph algorithm. Key findings from our extensive mathematical experiments on SynthLLM include: (1) SynthLLM generates synthetic data that reliably adheres to the rectified scaling law across various model sizes; (2) Performance improvements plateau near 300B tokens; and (3) Larger models approach optimal performance with fewer training tokens. For instance, an 8B model peaks at 1T tokens, while a 3B model requires 4T. Moreover, comparisons with existing synthetic data generation and augmentation methods demonstrate that SynthLLM achieves superior performance and scalability. Our findings highlight synthetic data as a scalable and reliable alternative to organic pre-training corpora, offering a viable path toward continued improvement in model performance.
Figure 1: Scaling laws on different model sizes. The x axis denotes the number of training tokens. The y axis represents the models’ error rates on MATH. The green points represent the data sizes used to fit the scaling laws, while the red points are used to test the prediction performance of the fitted curves.
1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across a wide range of real-world applications, including chat assistant [39, 4] and code generation [9, 3]. Their success is fundamentally driven by the availability of vast amounts of high-quality data, which fuels model development at scale. As an essential resource, pre-training data provides the fundamental knowledge and generalization capability that underpin LLMs’ broad applicability [41, 7]. The importance of pre-training data is further reinforced by scaling laws [20, 18, 38], which suggest that larger models trained on more data tend to exhibit predictable and consistent performance improvements.
However, recent studies [48, 37, 44] have raised concerns that our supply of high-quality pre-training data is rapidly depleting. With the web-scraped textual corpora that power LLMs approaching saturation, the conventional scaling paradigm may face diminishing returns, potentially slowing down future progress. Rather than continuing to rely solely on next-token prediction over raw pre-training data, it is crucial to explore more principled methods to better harness the potential of existing corpora. One promising direction is the use of synthetic data [13, 27, 30, 35, 32, 1], which offers a way to amplify and extend the utility of existing human-generated data, potentially sustaining performance improvements even as natural data resources dwindle. This prompts a natural question: Are there scaling laws for synthetic data? While scaling laws for pre-training data are well-documented [20, 18], whether similar principles apply to synthetic data remains unclear. Can scaling synthetic datasets sustain performance gains, or do fundamental limitations arise? Understanding its scaling behavior is key to assessing its long-term viability as a solution to data scarcity.
To investigate the scaling laws of synthetic data, we aim to design a scalable approach capable of generating synthetic data at the scale. Conventional approaches to synthetic dataset curation depend heavily on limited human-annotated seed examples from target domains [50, 43, 53, 36, 46, 12, 23]. This dependency fundamentally constrains both the diversity and scalability of the resulting datasets. In contrast, pre-training corpus—being both vast and highly diverse—still remains an underutilized resource for scalable synthetic data generation. We develop SynthLLM, a scalable web-scale synthetic data generation method designed to systematically transform pre-training data into high-quality synthetic datasets.
SynthLLM operates through three distinct stages. Initially, our method autonomously identifies and filters high-quality web documents within a target domain (e.g., Mathematics). Subsequently, leveraging these high-quality reference documents, we generate large-scale, diverse questions (or prompts) by employing open-source LLMs through three complementary methods, each strategically designed to progressively enhance question diversity. In final stage, we produce corresponding answers (or responses) to these generated questions, again utilizing open-source LLMs. It is worth noting that in the second stage, previous approaches have typically employed either direct question extraction approaches [55, 56, 54] or implemented document back-translation [26] to generate questions. However, these methods demonstrate inherent limitations in scalability, as question generation is constrained either by the finite size of reference documents containing high-quality questions or by the necessity of training dedicated back-translation models. Our method goes beyond direct extraction by automatically extracting and randomly combining high-level concepts from multiple documents with a graph algorithm, while establishing grounding across reference documents. Detailed ablation studies highlight the superior performance and scalability gained by our methods.
We apply SynthLLM to the mathematical reasoning domain to investigate scaling laws in synthetic data. We observe the following: (1) Synthetic data consistently follow the rectified scaling law [29] across various model sizes, as illustrated in Figure 1. (2) Performance gains start to diminish once the amount of synthetic data exceeds approximately 300B tokens. (3) Larger models reach optimal performance more quickly compared to smaller ones. For instance, the 8B model requires only 1T tokens to achieve its best performance, whereas the 3B model needs 4T tokens.
These results underscore the critical importance of scaling synthetic data. Even with limited pre-training data, systematically expanding the synthetic dataset yields sustained and predictable performance gains, reinforcing its role as a scalable solution to reduce reliance on human-generated data while enhancing model capabilities.
2 Related Work
Synthetic Data Generation.
Synthetic data has emerged as a promising approach to augment real data, addressing challenges related to data scarcity and the high cost of data collection and annotation. It has been widely applied in pre-training, instruction following, and reasoning tasks, demonstrating notable improvements in model performance. In this work, we focus on synthetic data generation for the post-training phase. Existing synthetic data generation methods typically rely on a small set of high-quality human-annotated seed samples, leveraging LLMs to generate diverse augmentations through various techniques: 1) sampling seed instructions as few-shot examples to elicit LLMs to generate new instructions [50, 19, 47, 23]; 2) leveraging LLMs to rephrase seed samples [53, 35, 12, 21, 34]. However, the scarcity of high-quality seed examples limits the scalability and diversity of generated data. To address this, Xu et al. [51] and Li et al. [24] propose generating synthetic data from scratch, leveraging either the inherent uncertainty of LLMs or a predefined knowledge taxonomy to guide the generation process. In contrast, pre-training data—being both vast and highly diverse—still remains an underutilized resource for scalable post-training data generation. Recent methods have explored extracting high-quality samples directly from web documents (similar to our Level-1 method) [55, 56, 54] or employing document backtranslation to generate questions [26]. Our SynthLLM aligns with this line of research and offers a more efficient approach to leveraging web documents for scaling diverse sample generation compared to aforementioned methods.
Scaling Law of LLMs.
Scaling laws provide a predictive framework for estimating model performance based on key factors such as model size and pre-training data size, and have been extensively studied [17, 42, 20, 18, 5]. These laws offer valuable insights into how performance scales with computational resources, enabling more informed decisions on the optimal allocation of compute for pre-training LLMs [20, 18]. Recently, more fine-grained scaling laws have been proposed, including data-constrained scaling [37], hyperparameter scaling [6], and model distillation [8]. For LLM fine-tuning, Hernandez et al. [16] investigate scaling laws in the transition from unsupervised pre-training to fine-tuning, highlighting the critical role of pre-training in altering the scaling behavior of the model. Lin et al. [29] propose a rectified scaling law tailored specifically for fine-tuning LLMs on downstream tasks. Our work extends the rectified scaling law [29] and is the first to investigate and empirically validate that scaling laws hold for fine-tuning language models with synthetic data.
3 Scaling Law of Synthetic Data
Kaplan et al. [20] demonstrates that the performance of LLMs exhibits a power-law relationship with respect to both model size and dataset size, primarily focusing on the pre-training phase utilizing organic data. In this section, we investigate the scaling properties of synthetic data. Specifically, we aim to determine whether a scaling law exists for synthetic data that enables accurate prediction of the performance of LLMs trained on such data. Details of the approach employed for synthetic data generation (i.e., SynthLLM) will be presented in Section 4, which is a method for curating synthetic data from carefully filtered documents drawn from the pre-training corpus (comprised exclusively of organic raw data).
3.1 Scaling Laws for Organic Data
The scaling law for pre-training on organic data has been extensively studied in the literature [20, 18]. In particular, the predictive loss is a parametric function of the model parameter size $N$ and data size $D$ (i.e., number of tokens) following the power-law form [18]:
$L(N,D)=\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}}+\hat{E}$ (1)
This function above models the relationship between scaling loss $L(N,D)$ (usually measured on a validation set), the number of model parameters $N$, and the number of training tokens $D$, based on the classical risk decomposition, where the final loss can be decomposed into multiple terms, each corresponding to a different source of error. $L(N,D)$ is the final scaling loss. $N$ represents parameter count of the LLM, and $D$ denotes the number of training tokens. $E$, the irreducible term, reflects the fundamental loss inherent in an ideal generative process operating over the natural text distribution. $\alpha$ and $\beta$ are decay exponents that determine how the loss decreases as $N$ and $D$ increase. $A$ and $B$ regulate the rate of loss convergence and are usually influenced by the model architecture and the characteristics of the data distribution. The coefficients $\alpha,\beta,A,B,$ and $E$ are learnable parameters that are determined by fitting a scaling curve.
Hernandez et al. [16] investigate scaling laws for transfer (i.e., the transition from unsupervised pre-training to fine-tuning). Their findings indicate that fine-tuning on pre-trained models exhibits more data efficient scaling behavior compared to training models from scratch on downstream tasks, emphasizing the crucial influence of pre-training on scaling dynamics. Extending this line of inquiry, Lin et al. [29] analyze scaling laws during fine-tuning of LLMs and propose Rectified Scaling Law:
$L(D)=\frac{B}{D_{l}+D^{\beta}}+E$ (2)
In the following, we illustrate how the rectified scaling law is obtained from the vanilla scaling law in Eq. (1).
Compared to the original formulation in Eq. (1), we first remove the term associated with model parameters, as the LLM remains fixed during fine-tuning. Consequently, we derive the following marginalized form of Eq. (1):
$L(D)=\frac{B}{D^{\beta}}+E$ (3)
Here, $L(D)$ represents the validation metric used to evaluate the LLM’s performance on the downstream task. The irreducible term $E$ empirically represents the optimal achievable model performance as the data size approaches infinity. This term depends not only on inherent loss of the data distribution but also on the capacity limitations of the model itself.
Compared to the formulation in Eq. (3), the Rectified Scaling Law introduces an additional term, $D_{l}$, which represents the pre-learned data size. This term quantifies the latent knowledge relevant to the downstream task that the model has already acquired during pre-training. Fine-tuning on a pretrained LLM often yields better performance than training from scratch on a downstream task. The inclusion of $D_{l}$ reflects this benefit. Consequently, we obtain the final formulation in Eq. (2). The term $\frac{B}{D_{l}}$ essentially defines the initial performance of the pre-trained LLM. Lin et al. [29] show that this modification offers a more precise representation of the impact of pre-training on the model performance in downstream tasks during fine-tuning.
3.2 Scaling Laws for Synthetic Data
Prior research [20, 18, 16, 29] has primarily focused on investigating the scaling properties of organic, real-world data and their impact on LLM performance. To our knowledge, this work is the first to systematically investigate and empirically validate that analogous scaling laws hold for synthetic data. We employ SynthLLM (Section 4) to generate data from a carefully filtered subset of the pre-training corpus, ensuring both scale and diversity (see Section 5.3 for details). Using the full synthetic dataset from the three methods in SynthLLM, we conduct continued training on LLMs of varying sizes (Llama-3.2-1B, Llama-3.2-3B, and Llama-3.1-8B) with progressively larger subsets (100K, 200K, 400K, 800K, 1.6M, 3.2M, and 7.4M examples). We selected the mathematical reasoning domain for our experiments due to its well-established evaluation protocols and metrics. Specifically, we report the error rate on the MATH dataset [15], which covers a wide range of mathematical subjects and incorporates varied difficulty levels (ranging from Level-1 to level 5), making it more comprehensive than alternative datasets. Note that we report error rate rather than accuracy to maintain consistency with previous scaling law literature [20, 18]. It is worth noting that both lower loss and lower error rate indicate superior performance in evaluations. See more details of our synthetic datasets and experiments in Section 5.1.
Overall Trend.
As shown in Figure 1, the synthetic data generated by SynthLLM adheres to the rectified scaling law (Equation 2). The $x$-axis represents the synthetic dataset size, measured in number of tokens, while the $y$-axis depicts the error rate on the MATH dataset. Both axes employ logarithmic scaling, consistent with previous work [20]. The green points denote the data sizes utilized to fit the scaling laws, whereas the red points serve to evaluate the predictive accuracy of the fitted curves. Our results demonstrate that synthetic data generated by SynthLLM consistently adheres to the rectified scaling law across various model sizes.
In addition to the rectified scaling law (Equation 2), we also attempted to fit our results with the power scaling law (Equation 3). Our analysis reveals that the power law fails to capture the scaling trend effectively (see Section 5.4). We hypothesize that this phenomenon may be attributed to the base models used in our experiments having already acquired certain mathematical problem-solving capabilities during their pre-training phase.
Model Size Matters.
The decay exponent $\beta$ in Eq. (2) quantifies the rate at which the error rate decreases as dataset size increases. A larger value of $\beta$ corresponds to a more rapid reduction in error rate, given the same amount of synthetic tokens. As in Figure 1, in comparison to its large counterparts (3B and 8B models), the 1B model exhibits a smaller $\beta$ value ($0.34$), suggesting that increasing the volume of synthetic data yields more modest performance improvements for this smaller model. The ratio $\frac{B}{D_{l}}$ demonstrates a strong correlation with the initial capacity of models, wherein a smaller value (smaller error rate) indicates greater initial model capability. The 7B model exhibits the smallest ratio value ($56.3$),
Model 10B tokens 50B tokens 250B tokens 300B tokens 1T tokens 4T tokens
3B $64.6$ $74.7$ $80.5$ $80.9$ $83.1$ $84.4$
8B $73.2$ $78.6$ $81.4$ $81.6$ $82.6$ $83.2$
Table 1: The predicted MATH accuracies (%) on 3B and 8B models when continuing to scale synthetic data based on our scaling law.
which may be attributed to its superior model capacity and more comprehensive pre-training data exposure. Furthermore, the 3B and 7B models demonstrate identical decay exponents ($\beta$), indicating that performance disparities stem predominantly from differences in initial knowledge acquisition and inherent model capacity rather than from variations in the efficacy of synthetic data scaling.
Figure 2: Scaling laws on different model sizes. The x-axis represents the number of training tokens. The y-axis shows the models’ error rates on MATH benchmark. Solid lines indicate fitted scaling curves for different model sizes. The dotted lines represent predicted final performances when synthetic data is further scaled, as extrapolated from these curves. More details are shown in Figure 1.
Predictable Performance.
For 3B and 8B models, we evaluated the predictive capability of our fitted scaling curves. As illustrated by the red squares in Figure 1, these curves accurately forecast error rates on the MATH benchmark for both models ($40.0\%$ and $28.7\%$, respectively), when synthetic dataset is expanded to 7.4M samples (4.5B tokens). These error rates correspond to accuracies of $60.0\%$ and $71.3\%$. This highlights the critical role of scaling synthetic data, demonstrating that even as pre-training data nears saturation, systematically expanding the synthetic dataset continues to yield sustained and predictable performance gains. Leveraging our established scaling curves, we can extrapolate the future performance of models with continued expansion of synthetic data beyond current volumes. Table 1 presents these projections for the 3B and 8B models. The results indicate that performance improvements begin to diminish after approximately 300B synthetic tokens. We estimate that the 3B model would require approximately 4T synthetic data tokens to approach its performance ceiling, while the 8B model would need approximately 1T synthetic tokens (also see Figure 2).
4 SynthLLM: Web-Scale Synthetic Data Generation
In this section, we present details of SynthLLM for synthetic data generation. For specific target domain (e.g., mathematical reasoning), we begin by filtering high-quality and domain-specific reference documents from Fineweb-Edu [40], an open repository of web data (Section 4.1). Next, using these filtered reference documents, we construct large-scale, diverse questions (or prompts) by prompting open-source LLMs through three distinct methods, each designed to progressively increase question diversity (Section 4.2). Finally, we produce corresponding answers (or responses) for the generated questions, again employing open-source LLMs (Section 4.3).
4.1 Reference Documents Filtering
Figure 3: Overview of reference documents filtering. The whole process contains two classifiers: 1) cold-start classifier trained on generated positive samples (documents) and randomly sampled negative examples; 2) fine-grained classifier trained on previously filtered documents and additional randomly sampled documents with extra ratings. We train fine-grained classifier for 2 iterations.
Cold-start Domain Classifier.
For a given target domain, we train a binary classifier to filter web documents from Fineweb-Edu. While negative examples (documents) can be randomly sampled, obtaining sufficient positive examples (documents) for each domain is more challenging. To address this, we use synthetic documents as positive examples. Specifically, we generate synthetic domain-specific documents based on syllabi from GLAN [24], which cover 126 distinct disciplines. Each syllabus is produced by GPT-4 "(almost) from scratch", with further details available in [24]. As illustrated in Figure 3, we first use these syllabi to prompt a large language model (LLM) to generate synthetic documents representative of the target domain (e.g., blog posts, exam papers, and textbooks). Once we obtain the synthetic positive examples and randomly sampled negative examples from Fineweb-Edu, we train a cold-start domain classifier. Applying this classifier yields an initial set of domain-relevant documents, $\mathcal{D}^{0}$, from Fineweb-Edu.
Fine-grained Domain Classifier.
The initial reference document set, $\mathcal{D}^{0}$, is typically small. To expand it, we iteratively classify Fineweb-Edu documents based on previously identified references, $\mathcal{D}^{t-1}$. At each iteration $t$, we randomly sample an additional subset $\mathcal{D}^{t-1}_{R}$ from Fineweb-Edu and instruct GPT-4o to evaluate all documents in $\mathcal{D}^{t-1}\cup\mathcal{D}^{t-1}_{R}$ on a 1–10 scale, considering relevance to the target domain, clarity, and language quality. These ratings are then used to train a fine-grained classifier, which we apply to Fineweb-Edu to construct the next iteration’s reference set, $\mathcal{D}^{t}$. We retain documents with scores above a threshold of 6.5, as it yields satisfactory performance in initial experiments. Empirically, two iterations suffice to generate a high-quality and sufficiently large reference corpus.
Both the cold-start and fine-grained classifiers are implemented using random forests, trained on Fineweb-Edu documents represented as vector embeddings generated by StableLM-2-1.6B.
4.2 Document-Grounded Question Generation
After applying the filtering process described in Section 4.1, we obtain the filtered document set $\mathcal{D}$. We then generate questions through three distinct methods, each relying on either single-document or multi-document contexts, with the goal of progressively increasing question diversity. Let $\mathbf{M}^{Q}$ denote the LLM employed for question generation.
Level-1 Generator.
In the first method, we generate questions from a single reference document $d\in\mathcal{D}$ (where $\mathcal{D}$ is filtered document set). Because $d$ may already contain questions relevant to target domain, we first prompt the LLM $\mathbf{M}^{Q}$ to determine whether $d$ includes any existing questions. If such questions exist, we directly extract them. We then instruct $\mathbf{M}^{Q}$ to create additional questions inspired by the content of $d$. Finally, $\mathbf{M}^{Q}$ annotates each question with a special tag, either $<$is_original$>$ or $<$newly_created$>$, to indicate whether the question is an original question from $d$. The detailed prompt for $\mathbf{M}^{Q}$ is in Figure 8 of Appendix. Similar directly extraction-based synthesis method has been adopted in Yue et al. [55] Zhou et al. [56], and Yuan et al. [54].
However, the Level-1 approach described above may be limited in scalability, as the diversity and content of newly generated questions is constrained by the finite set of high-quality reference documents $\mathcal{D}$ containing questions.
Level-2 Generator.
To address the aforementioned limitation and more effectively leverage the reference document $d$, we draw inspiration from pedagogical principles. In conventional educational settings [11, 2], instructors typically begin by introducing fundamental knowledge and key concepts. They then progressively guide students toward a deeper understanding and application of these concepts through practice or exercises, thereby fostering problem-solving and critical thinking skills. Our Level-2 method mirrors this approach and implement it in two stages, as illustrated in Figure 4
Figure 4: Illustrations of our Document-Grounded Question Generation methods.
First, we prompt the model to act as an instructor, extracting key topics and concepts from the document $d$ for teaching. Then, we randomly sample these extracted topics and key concepts and combine them with $d$ to guide further question generation. Here, the model again acts as an instructor, creating exercises that target the sampled topics and concepts. Specifically, we prompt $\mathbf{M}^{Q}$ with a prompt $\mathbf{p}_{k}$ to extract related topics (e.g., "Curvature", "Central Limit Theorem"), and key concepts (e.g., "Arc length of a curve", "Binomial Approximation to Normal Distribution"). Formally, we compute $\mathbf{K}=\mathbf{M}^{Q}(\mathbf{p}_{k},d)$, where $\mathbf{K}$ represents the extracted meta-information described above. Note that the detailed prompt $\mathbf{p}_{k}$ is provided in Figure 11 of Appendix. Figure 12 and 13 illustrate examples of the extracted $\mathbf{K}$.
In the second stage, we focus on generating questions. To enhance the diversity of generated questions, we inject randomness into the generation process. Specifically, we first instruct $\mathbf{M}^{Q}$ to randomly select and combine concepts from $\mathbf{K}$, producing a subset $\mathbf{k}_{s}$. The final question is generated by $\mathbf{M}^{Q}$, drawing on the reference document $d$, and the selected subset $\mathbf{k}_{s}$. We aim for the generated questions to remain grounded in the reference document $d$ and strictly adhere to the selected subset $\mathbf{k}_{s}$. We denote the whole question generation process above as follows:
$\mathbf{x}=\mathbf{M}^{Q}((\mathbf{p}_{q},\mathcal{R}(\mathbf{K}),d)),\ \ \ \mathbf{K}=\mathbf{M}^{Q}((\mathbf{p}_{k},d)),$ (4)
where $\mathbf{k}_{s}=\mathcal{R}(\mathbf{K})$, $\mathcal{R}$ is random selection operator and $\mathbf{x}$ is the generated questions. For simplicity, we integrate subset selection of $\mathbf{k}_{s}$ and the generation of the final questions into a single prompt $\mathbf{p}_{q}$. The detailed prompt $\mathbf{p}_{q}$ is shown in Figure 9 of Appendix. As indicated in Equation (4), the final questions $\mathbf{x}$ rely on both the (random) concept combinations generated by $\mathcal{R}(\mathbf{K})$ and the document $d$. This design intuitively offers greater diversity than the Level-1 method, which depends solely on $d$. This brings significant advantages. High-quality question synthesis is no longer constrained by the limited existing questions within reference documents. Instead, by decomposing and recombining key concepts, our Level-2 approach more efficiently exploits limited reference documents, thereby enabling more scalable synthetic question generation. In Section 5.3, we validate the benefits of Level-2 through a series of detailed ablation studies.
The overall set of topics and key concepts $\mathbf{K}$ resembles the syllabus from GLAN [24]; however, a key distinction is that each $\mathbf{K}$ in our approach is explicitly grounded in a reference document $d$.
Level-3 Generator.
Although the Level-2 generator can generate diverse questions by combining different topics and key concepts within a single document, each document inherently covers only a limited set of material. To further promote diversity of generation questions and make the approach more scalable, the Level-3 generator extends the Level-2 generator by incorporating concepts from multiple documents and grounding across multiple documents. The process of Level-3 generator is illustrated in Figure 4.
Global Concept Graph Construction.
In the Level-2 Generator, we have extracted topics and key concepts (i.e., $\mathbf{K}$) from a single document. In this section, we extend this modeling to a global perspective by considering relationships among all high-level concepts. Specifically, we construct a graph $G$, whose nodes comprise all distinct topics $\mathbb{T}$ and key concepts (KCs) $\mathbb{C}$. The intuition behind this construction is that topics or key concepts frequently co-occurring within the same document likely indicate meaningful associations. Thus, we utilize the co-occurrence statistics between topics and KCs to define the edge weights in the graph. Consequently, we consider three types of edges: topic-topic edges, topic-KC edges, and KC-KC edges, which results in three distinct sub-graphs: the topic graph, the topic-KC graph, and the KC graph. An edge between two nodes, a topic or KC $\mathbf{u}$ and another topic or KC $\mathbf{v}$, is established whenever they appear together within the same document. The edge weight $w(\mathbf{u},\mathbf{v})$ is formally defined as:
$w(\mathbf{u},\mathbf{v})=\log\big{(}\text{freq}(\mathbf{u},\mathbf{v})+\epsilon\big{)}$ (5)
where $\text{freq}(\mathbf{u},\mathbf{v})$ denotes the raw co-occurrence count of nodes $\mathbf{u}$ and $\mathbf{v}$, and $\epsilon$ is a small positive constant introduced to ensure numerical stability.
Concept Combination Sampling.
Once the concept graph $G$ has been constructed, we proceed to sample concept combinations. These sampled combinations are then utilized to retrieve relevant grounding documents, ultimately facilitating the generation of new questions.
The sampling process begins with uniform random sampling from the nodes in the topic sub-graph $\mathbb{T}$. In implementation, we iterate through all topics in $\mathbb{T}$ over multiple epochs to achieve this. Once an initial topic is randomly selected, we perform a random walk of one to two steps within the topic sub-graph to identify additional related topics. The transition probability of moving from node $\mathbf{u}$ to node $\mathbf{v}$ during the random walk is computed using a softmax function:
$p_{\mathbf{u,v}}=\frac{\exp\left(w(\mathbf{u},\mathbf{v})\right)}{\sum_{\mathbf{v}^{\prime}\in\mathcal{N}(\mathbf{u})}\exp\left(w(\mathbf{u},\mathbf{v}^{\prime})\right)}$ (6)
where $\mathcal{N}(\mathbf{u})$ denotes the set of nodes adjacent to $\mathbf{u}$ in the topic sub-graph.
After obtaining the sampled topics, we next sample their corresponding key concepts. We first select an initial key concept by performing a one-step random walk on the topic-KC sub-graph, with transition probabilities computed using Equation (6) but applied specifically within this sub-graph. Subsequently, additional related key concepts are sampled through three to four steps of random walks on the KC sub-graph. This process results in a set of sampled topics and key concept $\mathbf{K}^{g}$.
Since the sampled or extracted high-level concepts essentially serve as a summary, they may omit certain details originally present in the reference documents. Thus, we search for reference documents closely aligned with our sampled concept set $\mathbf{K}^{g}$. Specifically, we calculate the Jaccard similarity between the concept set $\mathbf{K}^{g}$ and the concept sets of each document within the filtered seed document set $\mathcal{D}$. Subsequently, we select the two documents with the highest similarity scores as the final reference documents. Analogous to the procedure in the Level-2 Generator, we instruct $\mathbf{M}^{Q}$ to randomly select and combine concepts from $\mathbf{K}^{g}$, and then utilize the selected concepts as well as the two chosen documents as references for question generation. The detailed prompt is shown in Figure 10 of Appendix. In the Level-3 Generator, the combination of concepts go beyond a single document. In addition to grounding on multiple documents, by employing random sampling on a concept graph constructed from the entire filtered document set, the diversity of the generated questions is significantly increased, thereby broadening the scope and richness of the resulting question set.
4.3 Answer Generation
Using methods described in Section 4.2, we obtain a large set questions. Now we are ready to generation answers. We employ open-source LLMs such as Qwen2.5-Math-72B-Instruct as our answer generator $\mathbf{M}^{A}$. We have not yet incorporated an additional answer verification process, since preliminary experiments indicate that our synthetic data has already yielded satisfactory results. We do believe that dedicating effort to constructing and validating answers (e.g., majority voting and multi-agent debating) can further enhance the quality of our synthetic data at additional computational cost. We plan to leave the answer verification process in further work.
5 Experiments
5.1 Experimental Settings
Models.
For $\mathbf{M}^{Q}$, we utilize Mistral-Large-Instruct-2407 [45], as we believe that larger models inherently possess the essential internal knowledge required for effective concept extraction and question synthesis. For the answer generator, $\mathbf{M}^{A}$, we employ Qwen2.5-Math-72B [52]. For trained models, we adopt the Llama-3.2-1B, Llama-3.2-3B, and Llama-3.1-8B base models [31].
Training Settings.
We employ Supervised Fine-Tuning (SFT) to train the model. We adopt AdamW optimizer [33] and set learning rate to $1\times 10^{-5}$ with a batch size of $512$. We utilize the linear learning rate scheduler and train the model for 3 epochs.
Our Datasets.
[…]
Statistics of Synthetic Questions and Answers.
The median lengths of the questions generated by Level-1, Level-2, and Level-3 are $32$, $66$, and $80$, respectively, while the corresponding median lengths of the answers are $358$, $497$, and $545$.
Evaluation.
To evaluate the math reasoning ability of our model trained on synthetic data, we utilize the evaluation code from Qwen-Math [52] to conduct assessments on several mainstream benchmarks: MATH [15], GSM8K [10], OlympiadBench-Math (Olympiad) [14], CollegeMath (College) [43], Minerva Math (Minerva) [22], and Gaokao 2023 En (Gaokao) [28].
Other Baselines.
We compare our synthetic data with other high-quality synthetic datasets. MAmmoTH2 [55], NaturalReasoning [54], and JiuZhang3.0 [56] employ a direct extraction-based synthesis approach on pre-training documents, similar to our Level-1 method. OpenMathInstruct-2 [47] is a large-scale synthetic math dataset consisting of 14M samples, with GSM8K and MATH serving as the seed datasets. Additionally, we compare our dataset with NuminaMath [25], a high-quality dataset that includes both synthetically generated data and extracted question-answer pairs from exam papers and mathematics discussion forums.
Furthermore, we evaluate our Level-2 and Level-3 approaches against existing methods designed to enhance question diversity through augmentation. Specifically, we include two popular augmentation approaches: the rephrase augmentation approach [53, 35] and the persona augmentation approach [12, 21], both of which generate new questions based on the questions extracted using Level-1. For rephrase augmentation, we follow the same procedure as MetaMath [53], providing two-shot examples in the prompt to rephrase questions. For persona augmentation, we utilize the 200K personas released by the authors as prompts for question augmentation.
5.2 Main Results
In this section, we present main evaluation results of SynthLLM. We test our models on various mathematical reasoning benchmarks, which encompass a range of difficulty levels. The results are shown in Table 2. The values in parentheses represent the sample size. For compared baseline training datasets, we directly evaluate model checkpoints they have released. We adopt the entire dataset synthesized with different methods from our framework to train models and denote it as SynthLLM-3B or -8B. The full dataset has a sample size of 7.4M, and we also create a subset with 3.2M samples.
SynthLLM achieves superior performance compared to other synthetic datasets.
SynthLLM consistently outperforms other synthetic datasets across all benchmarks, demonstrating the effectiveness of our framework. OpenMathInstruct-2, a high-quality math reasoning dataset generated using the Llama3.1-405B model based on MATH and GSM8K, comprises 14M samples. Despite its significantly larger size, the model trained on smaller dataset (SynthLLM-8B (3.2M)) achieves comparable performance on its in-distribution test sets (MATH and GSM8K) while substantially surpassing OpenMathInstruct-2 on Out-of-Distribution (OOD) test sets. This highlights the superior generalization capabilities of SynthLLM. Compared to the corresponding Llama Instruct models across different sizes, our models consistently achieve superior performance. Furthermore, our models (SynthLLM-8B) match or even surpass much larger models such as NuminaMath-CoT-72B and Llama-3.1-70B-Instruct on benchmarks, further reinforcing the effectiveness of our approach.
Model GSM8K MATH Minerva Olympiad College Gaokao Average
Llama-3.2-1B-Instruct $47.2$ $28.0$ $5.9$ $5.5$ $18.8$ $25.2$ $21.8$
SynthLLM-1B (3.2M) $45.7$ $32.9$ $6.6$ $7.6$ $27.3$ $28.8$ $24.8$
SynthLLM-1B (7.4M) 50.4 37.4 8.1 8.3 31.7 30.6 27.4
Llama-3.2-3B-Instruct $78.0$ $47.5$ $17.3$ $14.8$ $31.8$ $37.7$ $37.9$
SynthLLM-3B (3.2M) $78.1$ $54.3$ $16.9$ $17.5$ $38.3$ $46.5$ $41.9$
SynthLLM-3B (7.4M) 80.7 60.0 18.8 21.9 42.3 50.9 45.8
Llama-3.1-8B-Instruct $84.2$ $48.9$ $25.7$ $13.2$ $32.1$ $43.4$ $41.2$
Llama-3.1-70B-Instruct 94.5 $66.1$ 34.2 $29.6$ $41.4$ $56.6$ $54.2$
NuminaMath-CoT-7B $75.4$ $55.2$ $19.1$ $19.9$ $36.9$ $47.5$ $42.3$
NuminaMath-CoT-72B $90.8$ $66.7$ $25.0$ $32.6$ $39.7$ $54.0$ $51.5$
MAmmoTH2-Plus-8B (10M) $78.4$ $41.9$ $10.7$ $11.3$ $16.1$ $31.9$ $31.7$
JiuZhang3.0-8B (6M) $88.7$ $51.2$ $21.7$ $18.8$ $37.5$ $43.4$ $43.6$
NaturalReasoning-8B (2.8M)* - $55.6$
OpenMathInstruct-2-8B (14M) $91.1$ $67.5$ $22.5$ $27.7$ $39.2$ $53.5$ $50.3$
SynthLLM-8B (3.2M) $88.4$ $66.1$ $25.4$ $30.2$ $44.3$ $56.9$ $51.9$
SynthLLM-8B (7.4M) $92.1$ 71.3 $26.5$ 33.0 45.3 61.0 54.9
Table 2: Performance of models of different sizes on various math benchmarks. All metrics are reported as percentages (%). We evaluate three model sizes, with the best result highlighted in bolded. The number in () represents training sample number. NaturalReasoning-8B: Results are directly referenced from the original paper, as the full dataset and model have not been released.
More importantly, for each reference document, we currently generate only five questions per document for Level-1 and Level-2, and three questions for Level-3, which is far from reaching the full potential of our method’s document utilization (as shown in Figure 6). We can further enhance performance on benchmarks simply by increasing the number of synthesized questions per document.
5.3 Ablation Studies about Our SynthLLM Framework
In this section, we present detailed ablation studies about our Level-2 and Level-3 methods, demonstrating their advantages in scalability and diversity.
Model GSM8K MATH Minerva Olympiad College Gaokao Average
Llama-3.2-1B-Instruct 47.2 $28.0$ $5.9$ $5.5$ $18.8$ $25.2$ $21.8$
Level-1-1B (2.3M) $38.9$ $28.5$ $6.6$ $5.3$ $\mathbf{26.5}$ 27.0 $22.2$
Level-2-1B (2.6M) $40.1$ $27.2$ $4.8$ $5.5$ $23.3$ $24.9$ $21.0$
Level-3-1B (2.6M) $42.6$ 30.1 9.2 7.1 $25.0$ $26.2$ 23.4
Llama-3.2-3B-Instruct 78.0 $47.5$ $17.3$ $14.8$ $31.8$ $37.7$ $37.9$
Level-1-3B (2.3M) $71.8$ $46.8$ $16.2$ $13.6$ $36.3$ $42.3$ $37.8$
Level-2-3B (2.6M) $74.1$ 49.9 $17.2$ $15.3$ 37.4 43.4 $39.6$
Level-3-3B (2.6M) $77.0$ $49.1$ 18.0 17.2 37.4 $42.9$ 40.3
Table 3: Performance of 1B and 3B models on various math benchmarks. The number in ()()( ) represents the sample size. The best result is in bolded. All metrics are measured in percentage (%).
Comparison of Level-2 and Level-3 with directly extraction-based synthesis method.
We compare the directly extraction-based synthesis method (Level-1) with our Level-2 and Level-3 methods across models of varying sizes. The results are shown in Table 3. We observe that, on the 3B model, Level-2 and Level-3 significantly outperform Level-1 and instruct version across benchmarks. On the 1B model, Level-3 surpasses Level-1 by a large margin on most benchmarks, with the exception of College and Gaokao. Although each document is sampled with 5 questions, the sample size for Level-1 (2.3M) is significantly lower than that of Level-2 (2.6M).
Figure 5: Histogram of question similarity within the same document. The x-axis represents the cosine similarity between question pairs from the same document, while the y-axis denotes the frequency of each bin.
This is because Level-1 is based on direct extraction, which leads to a substantial amount of duplicate questions being generated. After removing duplicates, the sample size decreases considerably. This also highlights the scalability limitations of direct extraction-based synthesis, as it is inherently constrained by the availability of high-quality reference documents containing questions.
Diversity of Generated Questions. Compared to direct extraction-based synthesis, SynthLLM generates more diverse questions by decomposing and recombining knowledge concepts. To assess this, we compare Level-1 and Level-2 by measuring the similarity among generated questions within the same document. Specifically, we randomly select 2,000 documents and generate 50 questions per document for each method. We adopt all-MiniLM-L6-v2 [49] to obtain question embeddings and compute cosine similarity between each pair of questions within the same document. The histogram of question similarity is shown in Figure 5. We could clearly observe that questions generated by Level-2 exhibit lower similarity, indicating greater diversity among questions derived from the same document.
Figure 6: (a): The performances of Level-2, Level-3, and other augmentation methods on MATH benchmark. (b): the average performances across various benchmarks. The x axis denotes the sample number. The y axis represents the accuracy.
Comparison with other data augmentation methods on limited reference documents.
We compare our Level-2 and Level-3 methods with mainstream augmentation techniques, including rephrase [53] and persona augmentation [12]. We restrict the number of reference documents to a very small set and evaluate different methods for scaling high-quality question generation from a constrained document pool. The effectiveness of each method will be measured by its ability to consistently improve model performance on downstream tasks as the dataset expands. To achieve this goal, We randomly select 2,000 documents and generate 25, 50, 75, …, 150 questions per document—a significantly large number—using both Level-2 and Level-3 methods. This results in datasets of 50K, 100K, 150K, …, 300K questions for each method. For rephrase and persona augmentation methods, we first use Level-1 to extract existing questions from documents, yielding approximately 6,500 questions as seed dataset. We then expand this dataset to 300K questions using these two augmentation techniques. We train these datasets on Llama-3.1-8B model, with results presented in Figure 6. Figure 6 (a) illustrates MATH accuracy of different methods, while Figure 6 (b) shows average accuracy across benchmarks. Level-2 and Level-3 achieve significantly better performance on math benchmarks compared to augmentation methods. More importantly, as the number of generated questions per document increases to 150, our methods continue to show consistent improvements. In contrast, the performance of the rephrase and persona augmentation methods has already saturated as the dataset size expands. This demonstrates that compared with existing augmentation methods, our knowledge-guided methods more efficiently exploit limited reference documents by decomposing and recombining knowledge concepts, thereby enabling more scalable high-quality synthetic question generation. In previous experiments, we generated only 5 questions for Level-2 and 3 questions for Level-3. The results presented here indicate that our method has not yet reached its full potential in document utilization.
Figure 7: The results of fitting scaling curves using different formulations on the Llama-3.2-3B model. The x axis denotes the number of training tokens. The y axis represents the models’ error rates on MATH benchmark. The green points represent the data sizes used to fit the scaling laws, while the red points are used to test the prediction performance of the fitted curves.
5.4 Ablation Study on Scaling Law Form
We compare the rectified form Eq. (2) and marginal form Eq. (3) in terms of their effectiveness in fitting the scaling curve. The results of fitting scaling curves using these two formulations on the Llama-3.2-3B model are presented in Figure 7. The black dotted line represents the scaling curve fitted using the marginal form (Eq.(3)), while the green solid line corresponds to the fit using the rectified form (Eq.(2)). The green points indicate the data sizes used for fitting the scaling laws, whereas the red points are used to evaluate the predictive performance of the fitted curves. From the results, we observe that the marginal form fails to accurately fit the scaling curve and struggles to predict performance when scaling up the synthetic dataset. In contrast, the rectified form effectively captures the scaling law while also demonstrating strong predictive accuracy when scaling up the synthetic dataset. This underscores the importance of introducing the pre-learned size $D_{l}$ for scaling law of fine-tuning LLMs.
6 Conclusions and Future Work
In this work, we investigate the scaling laws of synthetic data and propose SynthLLM, a scalable web-scale data synthesis framework that systematically transforms pre-training data into high-quality synthetic datasets. Our findings show that synthetic data generated through SynthLLM consistently adheres to a rectified scaling law across various model sizes. By leveraging the fitted scaling law, we accurately predict model performance when the synthetic dataset size is doubled. These results suggest that systematically expanding synthetic datasets can continue to drive sustained and predictable performance improvements. Furthermore, our comparisons with other synthetic data generation and augmentation methods demonstrate that SynthLLM achieves superior performance and better scalability, providing a more effective approach for further enhancing model performance.
Our framework can be readily extended to other downstream domains, including code, physics and chemistry, and healthcare, expanding its applicability across diverse fields. Furthermore, SynthLLM is not limited to fine-tuning; we aim to explore its effectiveness in continued pre-training and even in the pre-training phase for synthetic data generation. Additionally, we will continue to refine and optimize our framework by developing more efficient strategies for leveraging pre-training data, with the goal of enhancing scaling efficiency of synthetic data.
Appendix A Used Prompts
In this section, we present prompts used in our SynthLLM framework.
Here is an article crawl from the web, which our classifier has identified as having significant educational value for students learning math. As a senior **math** instructor, your task is to create **challenging computation-based math questions**. These questions should be suitable for various contexts, such as homework assignments, exams, interview preparations, classroom activities, competitions, and tutoring sessions while enhancing students’ reasoning and critical-thinking skills. Ensure that questions are **non-redundant**, precise, and engaging. ### Guidelines for Creating Computation-based Questions: 1. **Assess Suitability**: If this article does not contain math-related content that can be used to generate engaging and solvable questions, please directly output "NOT SUITABLE for creating questions." 2. **Generate Questions**: If the article is suitable for creating math questions, generate **1 to 5** questions based on the richness and depth of the article content. For articles covering multiple topics, aim to generate more questions to ensure coverage:       - Each question must be solvable independently and should not rely on answers from previous questions.       - If a question closely resembles the original text, append "<original_question>" at the end of the question.       - If a question is newly created, append "<newly_created>" at the end of the question. 3. **Content Alignment**: Your math questions must exclusively draw from the content of the article, ensuring they are directly aligned with the concepts presented. 4. **Use Original Questions**: Use original questions from the article whenever possible. However, feel free to rephrase them for clarity and improved understanding. 5. **Create New Questions**: Attempt to formulate **newly diverse and challenging questions** that explore different aspects of the content presented in the article. 6. **Self-Contained Questions**: Ensure that each question is self-contained, meaning students do not need to read the article to answer them. 7. **Logical Consistency**: Verify that the questions are **logically sound and directly aligned with the mathematical principles** in the article. You MUST minimize the use of **sub-questions**, unless they are essential to the problem’s complexity. 8. **Clarity and Precision**:       - Use precise and unambiguous language.       - Write all mathematical expressions or formulas in LaTeX for clarity.       - Clearly state all assumptions or conditions.       - The answer should either be exact, or if not possible, then the question should clearly say the       answer is only expected to be approximately correct. #### Article {{ text }} #### Output Format For each question, provide the following information: - **Question Content**: The actual math question. - **School Level**: Specify the school level (e.g., <elementary>, <middle_school>, <high_school>, <college>, <grad_school>, <competition>). - **Originality Tag**: Append "<original_question>" for original questions from the article or"<newly_created>" for newly created questions. Example Output: <Q1> Question: Content Orig_tag:<original_question> Level:<high_school> </Q1> <Q2> Question: Content Orig_tag:<newly_created> Level:<college> </Q2>
Figure 8: Prompt for Level-1. The {{ text }} is the placeholder for inserting the reference document.
As a senior **math** instructor, your task is to create **diverse and challenging computation-based math questions**. These questions should demonstrate the application of the provided topics and key concepts while enhancing students’ reasoning and critical-thinking skills. Ensure that questions are **non-redundant**, precise, and engaging. ### Guidelines for Creating Diverse and Challenging Computation-based Questions: 1. **Concept Selection**:       - Randomly select **up to 2-3 distinct key concepts** from the provided list for each question.       - Ensure **broad coverage** of the provided concepts across the generated questions, avoiding over-reliance       on a limited subset of concepts.       - Avoid repeating the same **concept combinations** or **computational approach** across questions. 2. **Diversity and Challenge**:       - Ensure that each question explores **different combinations of key concepts** and is **sufficiently       challenging** (e.g., requiring multi-step computations, integrating real-world scenarios, involving abstract or       advanced reasoning.). 3. **Clarity and Precision**:       - Verify that the questions are **logically sound**.       - Use precise and unambiguous language.       - Write all mathematical expressions or formulas in LaTeX for clarity.       - Clearly state all assumptions or conditions. 4. **Reference Material**:       - Use the provided **reference article** as a source of inspiration for generating **unique, diverse, and       challenging questions**.       - The reference material is intended to:         - Supplement the concept list by introducing **novel perspectives**, **contexts**, or **applications**.         - Help create questions that are **more complex, realistic, or uncommon** in traditional teaching scenarios.         - Serve as a resource to craft **real-world scenarios** or **abstract extensions** beyond the given concepts. 5. **Output Diversity**:       - Create between **1 to 5 questions**.       - Ensure each question is unique in **structure**, **approach**, and **concept usage**.       - Minimize the use of **sub-questions**, unless they are essential to the problem’s complexity.       - The answer should either be exact, or if not possible, then the question should clearly say the answer is only       expected to be approximately correct. ### Inputs: - **Article**:       {{ text }} - **Concept List**:       {{ concept }} #### Output Format: <Q1> Selected Concepts: [Only insert 2-3 concepts here] Question: [Only insert question here] </Q1> <Q2> Selected Concepts: [Only insert 2-3 concepts here] Question: [Only insert question here] </Q2>
Figure 9: Prompt for Level-2. The {{ text }} and {{ concept }} are the placeholders for inserting the reference document and meta-information (e.g., topics, key concepts) extracted from document.
As a senior **math** instructor, your task is to create **diverse and challenging computation-based math questions** based on provided topics and knowledge points. These questions should demonstrate the application of the provided topics and key concepts while enhancing students’ reasoning and critical-thinking skills. Ensure that questions are **non-redundant**, precise, and engaging. You will be provided with a list of key mathematical concepts spanning various topics and two relevant reference materials. ### Guidelines for Creating Diverse and Challenging Computation-based Questions: 1. **Concept Selection**:       - Adhere to the Provided Topics: Ensure that each question aligns closely with the given topics.       - Incorporate Multiple Concepts about Different Topics: Each question should encompass **2 or 3 key       concepts about different math topics**.       - Ensure **broad coverage** of the provided concepts across the generated questions, avoiding       **over-reliance** on simple or common applications of concepts.       - Avoid repeating the same **concept combinations** or **computational approach** across questions. 2. **Diversity and Challenge**:       - Encourage **Cross-Topic Thinking**: By integrating concepts about different math topics, questions       will promote holistic understanding and application of mathematical principles.       - **Leverage the Two Reference Materials**: The combination of both reference materials provides       a **broader and more diverse context**, allowing for the creation of questions that explore a wider range       of scenarios and applications. Use this to generate questions that challenge students in both familiar and       novel contexts.       - Ensure questions explore **different perspectives** and **applications** of the key concepts. Ensure each       question is **sufficiently challenging** (e.g., requiring multi-step computations, integrating real-world       scenarios, involving abstract or advanced reasoning.). 3. **Clarity and Precision**:       - Use precise and unambiguous language.       - Write all mathematical expressions or formulas in LaTeX for clarity.       - Clearly state all assumptions or conditions. 4. **Reference Material**:       - Use the provided **reference articles about different topics** as sources of inspiration for       generating **unique, diverse, and challenging questions**.       - The combination of these two materials allows you to create questions with **more varied perspectives,       contexts, and applications**, which can help test students’ abilities to apply concepts in different situations.       - The reference material is intended to:         - Supplement the concept list by introducing **novel perspectives**, **contexts**, or **applications**.         - Help create questions that are **more complex, much harder, and uncommon** in traditional         teaching scenarios.         - Serve as a resource to craft **real-world scenarios** or **abstract extensions** beyond the         given concepts. 5. **Output Diversity**:       - Create between **1 to 3 questions**.       - Ensure each question is unique in **structure**, **approach**, and **concept usage**.       - Minimize the use of **sub-questions**, unless they are essential to the problem’s complexity.       - The answer should either be exact, or if not possible, then the question should clearly say the answer is       only expected to be approximately correct. ### Inputs: - **Article**:       {{ text }} - **Concept List**:       {{ concept }} #### Output Format: <Q1> Selected Concepts: [Only insert 2-3 concepts here] Question: [Only insert question here] </Q1> <Q2> Selected Concepts: [Only insert 2-3 concepts here] Question: [Only insert question here] </Q2>
Figure 10: Prompt for Level-3. The {{ text }} and {{ concept }} are the placeholders for inserting reference documents and meta-information sampled from knowledge graph.
Here is an article crawl from the web, which our classifier has identified as having significant educational value for students learning math. Your task is to analyze this article and extract educational materials, specifically focusing on topics and key concepts that can enhance students’ understanding of mathematics and improve their problem-solving skills. Pay special attention to uncommon but important mathematical concepts that are crucial for a deeper understanding. ## Tasks 1. **Determine Educational Level:**       - Identify the appropriate educational level for the article based on its content. Choose from the       following options:         - Primary School         - Middle School         - High School         - College         - Graduate School         - Competition         - Other 2. **Identify Subject Area:**       - Specify the primary subject area of mathematics to which the article belongs (e.g., Calculus,       Geometry, Algebra, etc.). 3. **Extract Topics and Key Concepts:**       - **Topics:**         - List **1 to 5** main topics covered in the article.         - Use terms commonly recognized in academia or industry.       - **Key Concepts:**         - For each identified topic, list **5 to 20** related key concepts.         - Ensure these concepts are clearly articulated using standard academic or industry terms. ## Guidelines: - **Terminology:** Use precise and widely recognized academic or industry terminology for subjects, topics, and key concepts to maintain consistency and clarity. - **Educational Level Selection:** If appropriate, restrict the educational level to one of the following: "Primary School", "Middle School", "High School", "College", "Graduate School", or "Competition" to ensure accurate categorization. ## Text {{ text }} ## Output Format <level>Educational Level</level> <subject>Subject Area</subject> <topic> Topics: 1. topic 1 2. topic 2 </topic> <key_concept> Key Concepts: 1. topic 1:       1.1. key concept       1.2. key concept        2. topic 2:       2.1. key concept       … … </key_concept> ## Output
Figure 11: Prompt for extracting topics and key concepts from the document. The {{ text }} is the placeholders for inserting reference document.
Appendix B Examples about Extracted Topics and Key Concepts from Reference Document
In this section, we demonstrate some examples about extracted topics and key concepts from reference document by using previous prompt shown in Figure 11. We adopt a relative large Language model, Mistral-Large-Instruct-2407. The examples are shown in Figure 12 and 13. We can observe that Mistral-Large-Instruct-2407 is capable of extracting well-structured topics and their corresponding key concepts from the document.
<level>High School</level> <subject>Trigonometry</subject> <topic> Topics: 1. Trigonometric Functions and Identities 2. Geometry on a Sphere 3. Applications of Trigonometry 4. Complex Numbers and Trigonometry 5. Derivations and Proofs in Trigonometry </topic> <key_concept> Key Concepts: 1. Trigonometric Functions and Identities:       1.1. Sine, Cosine, and Tangent Functions       1.2. Trigonometric Identities (e.g., Pythagorean, Co-function, Sum and Difference)       1.3. Law of Sines and Law of Cosines       1.4. Inverse Trigonometric Functions       1.5. Trigonometric Equations and Their Solutions 2. Geometry on a Sphere:       2.1. Latitude and Longitude       2.2. Circumference of a Circle on a Sphere       2.3. Equation of a Circle on a Sphere       2.4. Spherical Trigonometry       2.5. Haversine Formula for Distance Calculation 3. Applications of Trigonometry:       3.1. Circular Motion and Angular Velocity       3.2. Calculating Angles in Right Triangles       3.3. Real-world Problems Involving Trigonometry (e.g., Crossing a Canyon)       3.4. Computing Sine and Cosine without a Calculator       3.5. Practical Uses of Trigonometric Functions in Engineering and Physics 4. Complex Numbers and Trigonometry:       4.1. Euler’s Formula and De Moivre’s Theorem       4.2. Complex Numbers and Trigonometric Functions       4.3. Roots of Unity and Their Relationship to Trigonometry       4.4. Inverse Hyperbolic Functions (e.g., Arcsinh)       4.5. Complex Numbers in Trigonometric Identities 5. Derivations and Proofs in Trigonometry:       5.1. Derivation of Trigonometric Identities       5.2. Proof of the Law of Sines and Law of Cosines       5.3. Derivation of Sum and Difference Formulas       5.4. Proof of the Haversine Formula       5.5. Derivation of Heron’s Formula </key_concept>
Figure 12: An example about High School’s Trigonometry.
<level>College</level> <subject>Vector Calculus</subject> <topic> Topics: 1. Surface Integrals of Vector Fields 2. Oriented Surfaces 3. Unit Normal Vectors 4. Flux of a Vector Field 5. Parametric Surfaces </topic> <key_concept> Key Concepts: 1. Surface Integrals of Vector Fields:       1.1. Definition of surface integral of a vector field       1.2. Flux of a vector field across a surface       1.3. Application of surface integrals in fluid dynamics       1.4. Evaluation of surface integrals using parametric surfaces       1.5. Surface integrals over closed surfaces 2. Oriented Surfaces:       2.1. Definition of an oriented surface       2.2. Positive and negative orientations       2.3. Unit normal vectors and their role in orientation       2.4. Orientation conventions for closed surfaces       2.5. Impact of orientation on surface integrals 3. Unit Normal Vectors:       3.1. Definition and calculation of unit normal vectors       3.2. Gradient vector and its role in finding normal vectors       3.3. Normal vectors for surfaces given by z = f(x, y)       3.4. Normal vectors for parametric surfaces       3.5. Adjusting normal vectors to match desired orientation 4. Flux of a Vector Field:       4.1. Definition of flux       4.2. Physical interpretation of flux in fluid dynamics       4.3. Calculation of flux using surface integrals       4.4. Flux across closed surfaces       4.5. Application of flux in Gauss’s Law 5. Parametric Surfaces:       5.1. Definition and representation of parametric surfaces       5.2. Calculation of normal vectors for parametric surfaces       5.3. Evaluation of surface integrals using parametric surfaces       5.4. Parameterization of common surfaces (e.g., spheres, cylinders)       5.5. Conversion between parametric and non-parametric forms </key_concept>
Figure 13: An example about the College’s Vector Calculus.

### v3

@@
 Abstract
-Large language models (LLMs) achieve strong performance across diverse tasks, largely driven by high-quality web data used in pre-training.
-However, recent studies indicate this data source is rapidly depleting.
+Large language models (LLMs) achieve strong performance across diverse tasks, driven by high-quality web data used in pre-training.
+However, recent studies indicate web data is rapidly depleting.
 Synthetic data emerges as a promising alternative, but it remains unclear whether synthetic datasets exhibit predictable scalability comparable to raw pre-training data.
-In this work, we systematically investigate the scaling laws of synthetic data by introducing SynthLLM, a scalable framework that transforms pre-training corpora into diverse, high-quality synthetic datasets.
+In this work, we systematically investigate scaling laws of synthetic data by introducing SynthLLM, a scalable framework that transforms pre-training corpora into diverse, high-quality synthetic datasets.
 Our approach achieves this by automatically extracting and recombining high-level concepts across multiple documents using a graph algorithm.
-Key findings from our extensive mathematical experiments on SynthLLM include: (1) SynthLLM generates synthetic data that reliably adheres to the rectified scaling law across various model sizes; (2) Performance improvements plateau near 300B tokens; and (3) Larger models approach optimal performance with fewer training tokens.
+Key findings from our experiments with SynthLLM on math domain include: (1) SynthLLM generates synthetic data that reliably adheres to rectified scaling law across various model sizes; (2) Performance gains gradually diminish near 300B tokens; and (3) Larger models approach optimal performance with fewer training tokens.
 For instance, an 8B model peaks at 1T tokens, while a 3B model requires 4T.
 Moreover, comparisons with existing synthetic data generation and augmentation methods demonstrate that SynthLLM achieves superior performance and scalability.
-Our findings highlight synthetic data as a scalable and reliable alternative to organic pre-training corpora, offering a viable path toward continued improvement in model performance.
-Figure 1: Scaling laws on different model sizes.
-The x axis denotes the number of training tokens.
-The y axis represents the models’ error rates on MATH.
-The green points represent the data sizes used to fit the scaling laws, while the red points are used to test the prediction performance of the fitted curves.
+Our findings highlight synthetic data as a scalable and reliable alternative to raw pre-training data, offering a viable path toward continued improvement in model performance.
 1 Introduction
-Large language models (LLMs) have demonstrated remarkable capabilities across a wide range of real-world applications, including chat assistant [39, 4] and code generation [9, 3].
-Their success is fundamentally driven by the availability of vast amounts of high-quality data, which fuels model development at scale.
-As an essential resource, pre-training data provides the fundamental knowledge and generalization capability that underpin LLMs’ broad applicability [41, 7].
-The importance of pre-training data is further reinforced by scaling laws [20, 18, 38], which suggest that larger models trained on more data tend to exhibit predictable and consistent performance improvements.
-However, recent studies [48, 37, 44] have raised concerns that our supply of high-quality pre-training data is rapidly depleting.
-With the web-scraped textual corpora that power LLMs approaching saturation, the conventional scaling paradigm may face diminishing returns, potentially slowing down future progress.
-Rather than continuing to rely solely on next-token prediction over raw pre-training data, it is crucial to explore more principled methods to better harness the potential of existing corpora.
-One promising direction is the use of synthetic data [13, 27, 30, 35, 32, 1], which offers a way to amplify and extend the utility of existing human-generated data, potentially sustaining performance improvements even as natural data resources dwindle.
-This prompts a natural question: Are there scaling laws for synthetic data?
-While scaling laws for pre-training data are well-documented [20, 18], whether similar principles apply to synthetic data remains unclear.
-Can scaling synthetic datasets sustain performance gains, or do fundamental limitations arise?
-Understanding its scaling behavior is key to assessing its long-term viability as a solution to data scarcity.
-To investigate the scaling laws of synthetic data, we aim to design a scalable approach capable of generating synthetic data at the scale.
-Conventional approaches to synthetic dataset curation depend heavily on limited human-annotated seed examples from target domains [50, 43, 53, 36, 46, 12, 23].
-This dependency fundamentally constrains both the diversity and scalability of the resulting datasets.
-In contrast, pre-training corpus—being both vast and highly diverse—still remains an underutilized resource for scalable synthetic data generation.
-We develop SynthLLM, a scalable web-scale synthetic data generation method designed to systematically transform pre-training data into high-quality synthetic datasets.
+Large language models (LLMs) achieve remarkable capabilities, primarily driven by vast amounts of high-quality pre-training data, which provide the foundational knowledge and generalization capacity necessary for their broad applicability (Radford et al., 2019; Brown et al., 2020).
+This relationship is further reinforced by scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022; Muennighoff et al., 2023), indicating that increasing both model size and data volume typically yields predictable and consistent performance improvements.
+However, recent studies (Villalobos et al., 2024; Muennighoff et al., 2023) indicate that the supply of high-quality pre-training data is rapidly diminishing.
+As web-scraped textual corpora approach saturation, conventional scaling methods may face diminishing returns, potentially slowing future progress.
+Therefore, it becomes crucial to explore more principled approaches that better harness the potential of existing corpora.
+Synthetic data (Gunasekar et al., 2023; Li et al., 2023; Liu et al., 2024; Maini et al., 2024; Long et al., 2024; Abdin et al., 2024) emerges as a promising alternative, providing a means to amplify and extend the utility of human-generated data.
+Yet, it remains unclear whether synthetic data follows predictable scaling laws akin to those established for natural pre-training data (Kaplan et al., 2020; Hoffmann et al., 2022).
+This raises a critical question: Can synthetic data sustain scalable performance improvements, or are there fundamental limitations?
+Understanding this is essential to assessing the viability of synthetic data as a long-term solution to data scarcity.
+To investigate the scaling laws of synthetic data, we aim to develop a scalable approach for generating synthetic data at the scale.
+Conventional synthetic dataset generation methods rely heavily on limited human-annotated seed examples from target domains (Wang et al., 2022; Tang et al., ; Yu et al., 2024; Mitra et al., 2023; Toshniwal et al., 2025; Ge et al., 2024; Li et al., 2024a), fundamentally restricting the diversity and scalability of the resulting datasets.
+In contrast, pre-training corpora—vast and inherently diverse—remain an underutilized resource for synthetic data generation.
+To address this gap, we introduce SynthLLM, a scalable framework designed to systematically transform pre-training data into high-quality synthetic datasets.
 SynthLLM operates through three distinct stages.
-Initially, our method autonomously identifies and filters high-quality web documents within a target domain (e.g., Mathematics).
-Subsequently, leveraging these high-quality reference documents, we generate large-scale, diverse questions (or prompts) by employing open-source LLMs through three complementary methods, each strategically designed to progressively enhance question diversity.
-In final stage, we produce corresponding answers (or responses) to these generated questions, again utilizing open-source LLMs.
-It is worth noting that in the second stage, previous approaches have typically employed either direct question extraction approaches [55, 56, 54] or implemented document back-translation [26] to generate questions.
-However, these methods demonstrate inherent limitations in scalability, as question generation is constrained either by the finite size of reference documents containing high-quality questions or by the necessity of training dedicated back-translation models.
-Our method goes beyond direct extraction by automatically extracting and randomly combining high-level concepts from multiple documents with a graph algorithm, while establishing grounding across reference documents.
+First, it autonomously identifies and filters high-quality web documents from the target domain (e.g., mathematics).
+Second, using these selected reference documents, it generates diverse, large-scale questions or prompts through open-source LLMs employing three complementary methods designed to progressively enhance diversity.
+Finally, the method produces corresponding answers to these questions, again utilizing open-source LLMs.
+Previous approaches typically used direct question extraction (Yue et al., 2024; Zhou et al., 2024; Yuan et al., 2025) or document back-translation (Li et al., 2024).
+However, these methods have inherent scalability limitations, constrained by either the finite number of reference documents containing quality questions or the necessity of dedicated back-translation model training.
+SynthLLM goes beyond direct extraction by automatically extracting and randomly combining high-level concepts from multiple documents with a graph algorithm, while establishing grounding across diverse reference documents.
 Detailed ablation studies highlight the superior performance and scalability gained by our methods.
-We apply SynthLLM to the mathematical reasoning domain to investigate scaling laws in synthetic data.
-We observe the following: (1) Synthetic data consistently follow the rectified scaling law [29] across various model sizes, as illustrated in Figure 1.
+We apply SynthLLM to math reasoning domain to study the scaling laws of synthetic data.
+We observe the following: (1) Synthetic data consistently follow rectified scaling law (Lin et al., 2024) across various model sizes, as illustrated in Figure 1.
 (2) Performance gains start to diminish once the amount of synthetic data exceeds approximately 300B tokens.
@@
 Even with limited pre-training data, systematically expanding the synthetic dataset yields sustained and predictable performance gains, reinforcing its role as a scalable solution to reduce reliance on human-generated data while enhancing model capabilities.
-2 Related Work
-Synthetic Data Generation.
-Synthetic data has emerged as a promising approach to augment real data, addressing challenges related to data scarcity and the high cost of data collection and annotation.
-It has been widely applied in pre-training, instruction following, and reasoning tasks, demonstrating notable improvements in model performance.
-In this work, we focus on synthetic data generation for the post-training phase.
-Existing synthetic data generation methods typically rely on a small set of high-quality human-annotated seed samples, leveraging LLMs to generate diverse augmentations through various techniques: 1) sampling seed instructions as few-shot examples to elicit LLMs to generate new instructions [50, 19, 47, 23]; 2) leveraging LLMs to rephrase seed samples [53, 35, 12, 21, 34].
-However, the scarcity of high-quality seed examples limits the scalability and diversity of generated data.
-To address this, Xu et al.
-[51] and Li et al.
-[24] propose generating synthetic data from scratch, leveraging either the inherent uncertainty of LLMs or a predefined knowledge taxonomy to guide the generation process.
-In contrast, pre-training data—being both vast and highly diverse—still remains an underutilized resource for scalable post-training data generation.
-Recent methods have explored extracting high-quality samples directly from web documents (similar to our Level-1 method) [55, 56, 54] or employing document backtranslation to generate questions [26].
-Our SynthLLM aligns with this line of research and offers a more efficient approach to leveraging web documents for scaling diverse sample generation compared to aforementioned methods.
-Scaling Law of LLMs.
-Scaling laws provide a predictive framework for estimating model performance based on key factors such as model size and pre-training data size, and have been extensively studied [17, 42, 20, 18, 5].
-These laws offer valuable insights into how performance scales with computational resources, enabling more informed decisions on the optimal allocation of compute for pre-training LLMs [20, 18].
-Recently, more fine-grained scaling laws have been proposed, including data-constrained scaling [37], hyperparameter scaling [6], and model distillation [8].
-For LLM fine-tuning, Hernandez et al.
-[16] investigate scaling laws in the transition from unsupervised pre-training to fine-tuning, highlighting the critical role of pre-training in altering the scaling behavior of the model.
-Lin et al.
-[29] propose a rectified scaling law tailored specifically for fine-tuning LLMs on downstream tasks.
-Our work extends the rectified scaling law [29] and is the first to investigate and empirically validate that scaling laws hold for fine-tuning language models with synthetic data.
 2 Scaling Law of Synthetic Data
-Kaplan et al.
-[20] demonstrates that the performance of LLMs exhibits a power-law relationship with respect to both model size and dataset size, primarily focusing on the pre-training phase utilizing organic data.
-In this section, we investigate the scaling properties of synthetic data.
-Specifically, we aim to determine whether a scaling law exists for synthetic data that enables accurate prediction of the performance of LLMs trained on such data.
-Details of the approach employed for synthetic data generation (i.e., SynthLLM) will be presented in Section 4, which is a method for curating synthetic data from carefully filtered documents drawn from the pre-training corpus (comprised exclusively of organic raw data).
+In this section, we first investigate and demonstrate the scaling properties of synthetic data.
+Details of the approach employed for synthetic data generation (i.e., SynthLLM) will be presented in Section 3.
 2.1 Scaling Laws for Organic Data
-The scaling law for pre-training on organic data has been extensively studied in the literature [20, 18].
-In particular, the predictive loss is a parametric function of the model parameter size $N$ and data size $D$ (i.e., number of tokens) following the power-law form [18]:
+The scaling law for pre-training on organic data has been extensively studied in the literature (Kaplan et al., 2020; Hoffmann et al., 2022).
+In particular, the predictive loss is a parametric function of the model parameter size $N$ and data size $D$ (i.e., number of tokens) following the power-law form (Hoffmann et al., 2022):
 $L(N,D)=\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}}+\hat{E}$ (1)
-This function above models the relationship between scaling loss $L(N,D)$ (usually measured on a validation set), the number of model parameters $N$, and the number of training tokens $D$, based on the classical risk decomposition, where the final loss can be decomposed into multiple terms, each corresponding to a different source of error.
-$L(N,D)$ is the final scaling loss.
-$N$ represents parameter count of the LLM, and $D$ denotes the number of training tokens.
-$E$, the irreducible term, reflects the fundamental loss inherent in an ideal generative process operating over the natural text distribution.
-$\alpha$ and $\beta$ are decay exponents that determine how the loss decreases as $N$ and $D$ increase.
-$A$ and $B$ regulate the rate of loss convergence and are usually influenced by the model architecture and the characteristics of the data distribution.
-The coefficients $\alpha,\beta,A,B,$ and $E$ are learnable parameters that are determined by fitting a scaling curve.
+The function above models the scaling loss L(N,D)L(N,D)—typically measured on a validation set—as a relationship between the number of model parameters $N$ and training tokens $D$, based on classical risk decomposition.
+This decomposition splits the final loss into multiple terms, each corresponding to a distinct source of error.
+Here, $N$ denotes the LLM parameter count, and $D$ represents the number of training tokens.
+The irreducible term $E$ reflects the minimal achievable loss inherent to the ideal generative process over natural text distributions.
+The decay exponents $\alpha$ and $\beta$ control how rapidly the loss diminishes as $N$ and $D$ increase, respectively.
+The coefficients $A$ and $B$, typically influenced by model architecture and data characteristics, regulate the rate of loss convergence.
+All parameters $\alpha,\beta,A,B,$ and $E$ are learned by fitting the scaling curve to empirical data.
 Hernandez et al.
-[16] investigate scaling laws for transfer (i.e., the transition from unsupervised pre-training to fine-tuning).
-Their findings indicate that fine-tuning on pre-trained models exhibits more data efficient scaling behavior compared to training models from scratch on downstream tasks, emphasizing the crucial influence of pre-training on scaling dynamics.
-Extending this line of inquiry, Lin et al.
-[29] analyze scaling laws during fine-tuning of LLMs and propose Rectified Scaling Law:
+(2021) explore scaling laws for transfer learning, specifically the transition from unsupervised pre-training to fine-tuning.
+Their results indicate that fine-tuning pretrained models is more data-efficient than training from scratch, highlighting the importance of pre-training in shaping scaling behavior.
+Building upon this, Lin et al.
+(2024) further analyze scaling laws in fine-tuning scenarios and introduce the Rectified Scaling Law:
 $L(D)=\frac{B}{D_{l}+D^{\beta}}+E$ (2)
-In the following, we illustrate how the rectified scaling law is obtained from the vanilla scaling law in Eq. (1).
+In the following, we illustrate how the rectified scaling law is obtained from the vanilla scaling law in Eq.
+(1).
 Compared to the original formulation in Eq.
@@
 $L(D)=\frac{B}{D^{\beta}}+E$ (3)
-Here, $L(D)$ represents the validation metric used to evaluate the LLM’s performance on the downstream task.
-The irreducible term $E$ empirically represents the optimal achievable model performance as the data size approaches infinity.
-This term depends not only on inherent loss of the data distribution but also on the capacity limitations of the model itself.
-Compared to the formulation in Eq.
-(3), the Rectified Scaling Law introduces an additional term, $D_{l}$, which represents the pre-learned data size.
-This term quantifies the latent knowledge relevant to the downstream task that the model has already acquired during pre-training.
-Fine-tuning on a pretrained LLM often yields better performance than training from scratch on a downstream task.
-The inclusion of $D_{l}$ reflects this benefit.
+Here, $L(D)$ denotes the validation metric for evaluating the LLM’s performance on downstream tasks.
+The irreducible term $E$ represents the optimal performance achievable by the model as data size approaches infinity, reflecting both inherent data-distribution loss and model capacity constraints.
+Compared to Eq.
+(3), the Rectified Scaling Law introduces an additional parameter, $D_{l}$, termed the pre-learned data size.
+This quantifies the latent knowledge relevant to the downstream task acquired during pre-training.
+Including $D_{l}$ captures the empirical observation that fine-tuning pretrained models yields better performance than training from scratch.
 Consequently, we obtain the final formulation in Eq.
 (2).
-The term $\frac{B}{D_{l}}$ essentially defines the initial performance of the pre-trained LLM.
+Specifically, the term $\frac{B}{D_{l}}$ characterizes the initial performance advantage provided by pre-training.
 Lin et al.
-[29] show that this modification offers a more precise representation of the impact of pre-training on the model performance in downstream tasks during fine-tuning.
+(2024) demonstrate that this refinement provides a more accurate depiction of how pre-training influences downstream task performance during fine-tuning.
 2.2 Scaling Laws for Synthetic Data
-Prior research [20, 18, 16, 29] has primarily focused on investigating the scaling properties of organic, real-world data and their impact on LLM performance.
-To our knowledge, this work is the first to systematically investigate and empirically validate that analogous scaling laws hold for synthetic data.
+Prior studies (Kaplan et al., 2020; Hoffmann et al., 2022; Hernandez et al., 2021; Lin et al., 2024) have primarily investigated scaling laws of organic data.
+To our knowledge, this is the first systematic exploration and empirical validation of scaling laws for synthetic data.
 We employ SynthLLM (Section 3) to generate data from a carefully filtered subset of the pre-training corpus, ensuring both scale and diversity (see Section 4.3 for details).
 Using the full synthetic dataset from the three methods in SynthLLM, we conduct continued training on LLMs of varying sizes (Llama-3.2-1B, Llama-3.2-3B, and Llama-3.1-8B) with progressively larger subsets (100K, 200K, 400K, 800K, 1.6M, 3.2M, and 7.4M examples).
-We selected the mathematical reasoning domain for our experiments due to its well-established evaluation protocols and metrics.
-Specifically, we report the error rate on the MATH dataset [15], which covers a wide range of mathematical subjects and incorporates varied difficulty levels (ranging from Level-1 to level 5), making it more comprehensive than alternative datasets.
-Note that we report error rate rather than accuracy to maintain consistency with previous scaling law literature [20, 18].
-It is worth noting that both lower loss and lower error rate indicate superior performance in evaluations.
-See more details of our synthetic datasets and experiments in Section 5.1.
+In this work, we main focus on the mathematical reasoning domain due to its well-established evaluation protocols and metrics.
+Additionally, we conduct preliminary experiments in the coding domain, and present detailed results in Section 4.4.
+Model performance is assessed by reporting error rates on the MATH dataset (Hendrycks et al., 2021), which covers a wide range of mathematical subjects and difficulty levels (1–5).
+We use error rate as the evaluation metric, where lower error rates indicate better performance.
+Further details on datasets and experiments are provided in Section 4.1.
+Figure 1: Scaling laws across different model sizes.
+The x-axis denotes the number of training tokens, while the y-axis shows error rates on the MATH benchmark.
+Green points indicate data sizes used for fitting the scaling curves, and red points assess the predictive accuracy of these fitted curves.
 Overall Trend.
-As shown in Figure 1, the synthetic data generated by SynthLLM adheres to the rectified scaling law (Equation 2).
+As shown in Figure 1, synthetic data generated by SynthLLM adheres to the rectified scaling law (Equation 2).
 The $x$-axis represents the synthetic dataset size, measured in number of tokens, while the $y$-axis depicts the error rate on the MATH dataset.
-Both axes employ logarithmic scaling, consistent with previous work [20].
+Both axes employ logarithmic scaling, consistent with previous work (Kaplan et al., 2020).
 The green points denote the data sizes utilized to fit the scaling laws, whereas the red points serve to evaluate the predictive accuracy of the fitted curves.
-Our results demonstrate that synthetic data generated by SynthLLM consistently adheres to the rectified scaling law across various model sizes.
-In addition to the rectified scaling law (Equation 2), we also attempted to fit our results with the power scaling law (Equation 3).
-Our analysis reveals that the power law fails to capture the scaling trend effectively (see Section 5.4).
-We hypothesize that this phenomenon may be attributed to the base models used in our experiments having already acquired certain mathematical problem-solving capabilities during their pre-training phase.
-Model Size Matters.
-The decay exponent $\beta$ in Eq.
-(2) quantifies the rate at which the error rate decreases as dataset size increases.
-A larger value of $\beta$ corresponds to a more rapid reduction in error rate, given the same amount of synthetic tokens.
-As in Figure 1, in comparison to its large counterparts (3B and 8B models), the 1B model exhibits a smaller $\beta$ value ($0.34$), suggesting that increasing the volume of synthetic data yields more modest performance improvements for this smaller model.
-The ratio $\frac{B}{D_{l}}$ demonstrates a strong correlation with the initial capacity of models, wherein a smaller value (smaller error rate) indicates greater initial model capability.
-The 7B model exhibits the smallest ratio value ($56.3$),
+We additionally attempted fitting our data using Equation 3, but found that it fails to accurately model our results (see Section C.2).
+We hypothesize this discrepancy arises because the base models employed in our experiments already possess certain mathematical problem-solving capabilities acquired during pre-training.
+Model Size Matters The decay exponent $\beta$ in Eq.
+(2) quantifies how quickly the error rate decreases as dataset size grows; a larger $\beta$ indicates a more rapid improvement given same synthetic token volume.
+As shown in Figure 1, 1B model exhibits a smaller $\beta$ ($0.34$) compared to larger models (3B and 8B), suggesting it benefits less from increased synthetic data.
+The ratio $\frac{B}{D_{l}}$ correlates strongly with initial model performance, where a smaller ratio implies higher initial capability.
+The 7B model achieves the lowest ratio ($56.3$), reflecting its superior model capacity and richer pre-training exposure.
+Besides, identical decay exponents ($\beta$) for 3B and 7B models indicate that performance differences primarily arise from disparities in initial knowledge and model capacity rather than synthetic data scaling efficacy.
+Predictable Performance We evaluated the predictive accuracy of our fitted scaling curves for the 3B and 8B models.
+As shown by the red squares in Figure 1, these curves reliably forecast the models’ error rates (40.0% for 3B, 28.7% for 8B) on the MATH benchmark when scaled to 4.5B synthetic tokens, corresponding to accuracies of 60.0% and 71.3%, respectively.
+These results underscore the effectiveness of scaling synthetic data: even as traditional pre-training data approaches saturation, systematic expansion of synthetic datasets continues to provide stable and predictable performance improvements.
+Using our scaling curves, we further extrapolate the future performance of the 3B and 8B models as the volume of synthetic data continues to grow.
+Table 1 presents these projections, indicating diminishing returns in performance improvements beyond approximately 300B synthetic tokens.
 Model 10B tokens 50B tokens 250B tokens 300B tokens 1T tokens 4T tokens
@@
 Table 1: The predicted MATH accuracies (%) on 3B and 8B models when continuing to scale synthetic data based on our scaling law.
-which may be attributed to its superior model capacity and more comprehensive pre-training data exposure.
-Furthermore, the 3B and 7B models demonstrate identical decay exponents ($\beta$), indicating that performance disparities stem predominantly from differences in initial knowledge acquisition and inherent model capacity rather than from variations in the efficacy of synthetic data scaling.
-Figure 2: Scaling laws on different model sizes.
-The x-axis represents the number of training tokens.
-The y-axis shows the models’ error rates on MATH benchmark.
-Solid lines indicate fitted scaling curves for different model sizes.
-The dotted lines represent predicted final performances when synthetic data is further scaled, as extrapolated from these curves.
-More details are shown in Figure 1.
-Predictable Performance.
-For 3B and 8B models, we evaluated the predictive capability of our fitted scaling curves.
-As illustrated by the red squares in Figure 1, these curves accurately forecast error rates on the MATH benchmark for both models ($40.0\%$ and $28.7\%$, respectively), when synthetic dataset is expanded to 7.4M samples (4.5B tokens).
-These error rates correspond to accuracies of $60.0\%$ and $71.3\%$.
-This highlights the critical role of scaling synthetic data, demonstrating that even as pre-training data nears saturation, systematically expanding the synthetic dataset continues to yield sustained and predictable performance gains.
-Leveraging our established scaling curves, we can extrapolate the future performance of models with continued expansion of synthetic data beyond current volumes.
-Table 1 presents these projections for the 3B and 8B models.
-The results indicate that performance improvements begin to diminish after approximately 300B synthetic tokens.
-We estimate that the 3B model would require approximately 4T synthetic data tokens to approach its performance ceiling, while the 8B model would need approximately 1T synthetic tokens (also see Figure 2).
+Specifically, we estimate the 3B model would require around 4T synthetic tokens to approach its performance ceiling, whereas 8B model would need about 1T tokens.
+Currently, SynthLLM generates only five questions per document from pre-training corpus, which is still far from fully leveraging the potential of each document (see Figure 5).
+We can easily continue scaling up our synthetic dataset to further approach the performance ceiling.
 3 SynthLLM: Web-Scale Synthetic Data Generation
-In this section, we present details of SynthLLM for synthetic data generation.
-For specific target domain (e.g., mathematical reasoning), we begin by filtering high-quality and domain-specific reference documents from Fineweb-Edu [40], an open repository of web data (Section 4.1).
-Next, using these filtered reference documents, we construct large-scale, diverse questions (or prompts) by prompting open-source LLMs through three distinct methods, each designed to progressively increase question diversity (Section 4.2).
-Finally, we produce corresponding answers (or responses) for the generated questions, again employing open-source LLMs (Section 4.3).
+In this section, we detail the SynthLLM framework for synthetic data generation.
+For a given target domain (e.g., mathematical reasoning), we first filter high-quality, domain-specific reference documents from Fineweb-Edu (Penedo et al., 2024), a publicly available web data repository.
+Using these filtered documents, we then generate large-scale, diverse questions (or prompts) by leveraging open-source LLMs through three distinct methods, each designed to progressively enhance question diversity.
+Finally, we generate corresponding answers to these questions, again utilizing open-source LLMs.
 3.1 Reference Documents Filtering
 Figure 2: Overview of reference documents filtering.
-The whole process contains two classifiers: 1) cold-start classifier trained on generated positive samples (documents) and randomly sampled negative examples; 2) fine-grained classifier trained on previously filtered documents and additional randomly sampled documents with extra ratings.
-We train fine-grained classifier for 2 iterations.
 Cold-start Domain Classifier.
 For a given target domain, we train a binary classifier to filter web documents from Fineweb-Edu.
-While negative examples (documents) can be randomly sampled, obtaining sufficient positive examples (documents) for each domain is more challenging.
+While negative examples can be randomly sampled, obtaining sufficient positive examples for each domain is more challenging.
 To address this, we use synthetic documents as positive examples.
-Specifically, we generate synthetic domain-specific documents based on syllabi from GLAN [24], which cover 126 distinct disciplines.
-Each syllabus is produced by GPT-4 "(almost) from scratch", with further details available in [24].
+Specifically, we generate synthetic domain-specific documents based on syllabi from GLAN (Li et al., 2024b), which cover 126 distinct disciplines.
+Each syllabus is produced by GPT-4 "(almost) from scratch", with further details available in Li et al.
+(2024b).
 As illustrated in Figure 2, we first use these syllabi to prompt a large language model (LLM) to generate synthetic documents representative of the target domain (e.g., blog posts, exam papers, and textbooks).
@@
 3.2 Document-Grounded Question Generation
-After applying the filtering process described in Section 4.1, we obtain the filtered document set $\mathcal{D}$.
-We then generate questions through three distinct methods, each relying on either single-document or multi-document contexts, with the goal of progressively increasing question diversity.
-Let $\mathbf{M}^{Q}$ denote the LLM employed for question generation.
+After applying the filtering process outlined in Section 3.1, we obtain the refined document set $\mathcal{D}$.
+To generate questions, we employ three distinct methods, utilizing either single-document or multi-document contexts to progressively enhance question diversity.
+Let $\mathbf{M}^{Q}$ denote the LLM used for question generation.
 Level-1 Generator.
-In the first method, we generate questions from a single reference document $d\in\mathcal{D}$ (where $\mathcal{D}$ is filtered document set).
-Because $d$ may already contain questions relevant to target domain, we first prompt the LLM $\mathbf{M}^{Q}$ to determine whether $d$ includes any existing questions.
-If such questions exist, we directly extract them.
-We then instruct $\mathbf{M}^{Q}$ to create additional questions inspired by the content of $d$.
-Finally, $\mathbf{M}^{Q}$ annotates each question with a special tag, either $<$is_original$>$ or $<$newly_created$>$, to indicate whether the question is an original question from $d$.
+In the first method, we generate questions from a single reference document $d\in\mathcal{D}$.
+Since $d$ may already contain questions relevant to the target domain, we first prompt the LLM $\mathbf{M}^{Q}$ to determine whether $d$ includes any existing questions.
+If such questions exist, we directly extract them and encourage $\mathbf{M}^{Q}$ to rephrase them freely for clarity and improved understanding.
+We also instruct $\mathbf{M}^{Q}$ to create additional questions inspired by the content of $d$.
+Finally, $\mathbf{M}^{Q}$ annotates each question with a special tag, either <<original_question>> or <<newly_created>>, to indicate whether the question is an original question from $d$.
 The detailed prompt for $\mathbf{M}^{Q}$ is in Figure 7 of Appendix.
 Similar directly extraction-based synthesis method has been adopted in Yue et al.
-[55] Zhou et al.
-[56], and Yuan et al.
-[54].
+(2024), Zhou et al.
+(2024), and Yuan et al.
+(2025).
 However, the Level-1 approach described above may be limited in scalability, as the diversity and content of newly generated questions is constrained by the finite set of high-quality reference documents $\mathcal{D}$ containing questions.
 Level-2 Generator.
-To address the aforementioned limitation and more effectively leverage the reference document $d$, we draw inspiration from pedagogical principles.
-In conventional educational settings [11, 2], instructors typically begin by introducing fundamental knowledge and key concepts.
-They then progressively guide students toward a deeper understanding and application of these concepts through practice or exercises, thereby fostering problem-solving and critical thinking skills.
-Our Level-2 method mirrors this approach and implement it in two stages, as illustrated in Figure 4
+To overcome limitations of previous approach and better utilize $d$, we draw inspiration from pedagogical principles.
+In conventional education (Council et al., 2005; Anderson & Krathwohl, 2001), instructors first introduce fundamental knowledge and key concepts before guiding students toward deeper understanding through exercises, fostering problem-solving and critical thinking skills.
+Our Level-2 method mirrors this approach and implement it in two stages, as illustrated in Figure 3.
 Figure 3: Illustrations of our Document-Grounded Question Generation methods.
-First, we prompt the model to act as an instructor, extracting key topics and concepts from the document $d$ for teaching.
-Then, we randomly sample these extracted topics and key concepts and combine them with $d$ to guide further question generation.
-Here, the model again acts as an instructor, creating exercises that target the sampled topics and concepts.
-Specifically, we prompt $\mathbf{M}^{Q}$ with a prompt $\mathbf{p}_{k}$ to extract related topics (e.g., "Curvature", "Central Limit Theorem"), and key concepts (e.g., "Arc length of a curve", "Binomial Approximation to Normal Distribution").
-Formally, we compute $\mathbf{K}=\mathbf{M}^{Q}(\mathbf{p}_{k},d)$, where $\mathbf{K}$ represents the extracted meta-information described above.
-Note that the detailed prompt $\mathbf{p}_{k}$ is provided in Figure 11 of Appendix.
-Figure 12 and 13 illustrate examples of the extracted $\mathbf{K}$.
-In the second stage, we focus on generating questions.
-To enhance the diversity of generated questions, we inject randomness into the generation process.
-Specifically, we first instruct $\mathbf{M}^{Q}$ to randomly select and combine concepts from $\mathbf{K}$, producing a subset $\mathbf{k}_{s}$.
-The final question is generated by $\mathbf{M}^{Q}$, drawing on the reference document $d$, and the selected subset $\mathbf{k}_{s}$.
+First, we prompt the model to act as an instructor, extracting key topics and concepts from $d$ to serve as the foundation for question generation.
+Specifically, we prompt $\mathbf{M}^{Q}$ with $\mathbf{p}_{k}$ to identify relevant topics (e.g., "Curvature", "Central Limit Theorem") and key concepts (e.g., "Arc length of a curve", "Binomial Approximation to Normal Distribution"), formally defined as: $\mathbf{K}=\mathbf{M}^{Q}(\mathbf{p}_{k},d)$, where $\mathbf{K}$ represents the extracted meta-information.
+The detailed prompt $\mathbf{p}_{k}$ is provided in Figure 10, while Figure 11 and 12 illustrate examples of extracted $\mathbf{K}$.
+In the second stage, we generate questions while introducing controlled randomness to enhance diversity.
+Specifically, we instruct $\mathbf{M}^{Q}$ to randomly select and combine concepts from $\mathbf{K}$, forming a subset $\mathbf{k}_{s}$.
+The final question is then generated by $\mathbf{M}^{Q}$, conditioned on both the reference document $d$ and $\mathbf{k}_{s}$.
 We aim for the generated questions to remain grounded in the reference document $d$ and strictly adhere to the selected subset $\mathbf{k}_{s}$.
-We denote the whole question generation process above as follows:
+The overall question generation process is formally defined as:
 $\mathbf{x}=\mathbf{M}^{Q}((\mathbf{p}_{q},\mathcal{R}(\mathbf{K}),d)),\ \ \ \mathbf{K}=\mathbf{M}^{Q}((\mathbf{p}_{k},d)),$ (4)
 where $\mathbf{k}_{s}=\mathcal{R}(\mathbf{K})$, $\mathcal{R}$ is random selection operator and $\mathbf{x}$ is the generated questions.
-For simplicity, we integrate subset selection of $\mathbf{k}_{s}$ and the generation of the final questions into a single prompt $\mathbf{p}_{q}$.
-The detailed prompt $\mathbf{p}_{q}$ is shown in Figure 9 of Appendix.
-As indicated in Equation (4), the final questions $\mathbf{x}$ rely on both the (random) concept combinations generated by $\mathcal{R}(\mathbf{K})$ and the document $d$.
-This design intuitively offers greater diversity than the Level-1 method, which depends solely on $d$.
+For simplicity, we integrate subset selection of $\mathbf{k}_{s}$ and the generation of the final questions into a single prompt $\mathbf{p}_{q}$, detailed in Figure 8 of the Appendix.
+As shown in Eq.
+(4), $\mathbf{x}$ rely on both (random) concept combinations generated by $\mathcal{R}(\mathbf{K})$ and $d$.
+This design intuitively offers greater diversity than Level-1 method, which depends solely on $d$.
 This brings significant advantages.
@@
 Instead, by decomposing and recombining key concepts, our Level-2 approach more efficiently exploits limited reference documents, thereby enabling more scalable synthetic question generation.
-In Section 5.3, we validate the benefits of Level-2 through a series of detailed ablation studies.
-The overall set of topics and key concepts $\mathbf{K}$ resembles the syllabus from GLAN [24]; however, a key distinction is that each $\mathbf{K}$ in our approach is explicitly grounded in a reference document $d$.
+In Section 4.3, we empirically validate these benefits through a series of ablation studies.
+The set of topics and key concepts $\mathbf{K}$ resembles the syllabus from GLAN (Li et al., 2024b), but unlike GLAN, each $\mathbf{K}$ in our approach is explicitly grounded in a reference document $d$.
 Level-3 Generator.
-Although the Level-2 generator can generate diverse questions by combining different topics and key concepts within a single document, each document inherently covers only a limited set of material.
-To further promote diversity of generation questions and make the approach more scalable, the Level-3 generator extends the Level-2 generator by incorporating concepts from multiple documents and grounding across multiple documents.
-The process of Level-3 generator is illustrated in Figure 4.
+Although Level-2 method enhances question diversity by combining different topics and key concepts within a single document, each document inherently covers only a limited scope of material.
+To further improve diversity and scalability, the Level-3 generator extends the Level-2 approach by incorporating concepts from multiple documents and establishing cross-document grounding.
+The process is illustrated in Figure 3.
 Global Concept Graph Construction.
-In the Level-2 Generator, we have extracted topics and key concepts (i.e., $\mathbf{K}$) from a single document.
-In this section, we extend this modeling to a global perspective by considering relationships among all high-level concepts.
-Specifically, we construct a graph $G$, whose nodes comprise all distinct topics $\mathbb{T}$ and key concepts (KCs) $\mathbb{C}$.
-The intuition behind this construction is that topics or key concepts frequently co-occurring within the same document likely indicate meaningful associations.
-Thus, we utilize the co-occurrence statistics between topics and KCs to define the edge weights in the graph.
-Consequently, we consider three types of edges: topic-topic edges, topic-KC edges, and KC-KC edges, which results in three distinct sub-graphs: the topic graph, the topic-KC graph, and the KC graph.
-An edge between two nodes, a topic or KC $\mathbf{u}$ and another topic or KC $\mathbf{v}$, is established whenever they appear together within the same document.
-The edge weight $w(\mathbf{u},\mathbf{v})$ is formally defined as:
+In the Level-2 method, topics and key concepts (i.e., $\mathbf{K}$) are extracted from a single document.
+Here, we extend this modeling to a global level by capturing relationships among all high-level concepts across documents.
+Specifically, we construct a graph $G$, where nodes represent all distinct topics $\mathbb{T}$ and key concepts (KCs) $\mathbb{C}$.
+The intuition behind this construction is that topics and key concepts frequently co-occurring within the same document likely indicate meaningful associations.
+Thus, we define edge weights based on co-occurrence statistics.
+The graph consists of three types of edges: topic-topic edges, topic-KC edges, and KC-KC edges, forming three subgraphs: the topic graph, the topic-KC graph, and the KC graph.
+An edge is established between two nodes, $\mathbf{u}$ and $\mathbf{v}$, whenever they co-occur in the same document.
+The edge weight $w(\mathbf{u},\mathbf{v})$ is defined as:
 $w(\mathbf{u},\mathbf{v})=\log\big(\text{freq}(\mathbf{u},\mathbf{v})+\epsilon\big)$ (5)
-where $\text{freq}(\mathbf{u},\mathbf{v})$ denotes the raw co-occurrence count of nodes $\mathbf{u}$ and $\mathbf{v}$, and $\epsilon$ is a small positive constant introduced to ensure numerical stability.
+where $\text{freq}(\mathbf{u},\mathbf{v})$ denotes the raw co-occurrence count of $\mathbf{u}$ and $\mathbf{v}$, and $\epsilon$ is a small positive constant for numerical stability.
 Concept Combination Sampling.
-Once the concept graph $G$ has been constructed, we proceed to sample concept combinations.
-These sampled combinations are then utilized to retrieve relevant grounding documents, ultimately facilitating the generation of new questions.
 The sampling process begins with uniform random sampling from the nodes in the topic sub-graph $\mathbb{T}$.
-In implementation, we iterate through all topics in $\mathbb{T}$ over multiple epochs to achieve this.
-Once an initial topic is randomly selected, we perform a random walk of one to two steps within the topic sub-graph to identify additional related topics.
-The transition probability of moving from node $\mathbf{u}$ to node $\mathbf{v}$ during the random walk is computed using a softmax function:
+We iterate through all topics in $\mathbb{T}$ over multiple epochs.
+After selecting an initial topic at random, we perform a random walk of one to two steps within the topic sub-graph to identify additional related topics.
+The transition probability from node $\mathbf{u}$ to node $\mathbf{v}$ is computed as:
 $p_{\mathbf{u,v}}=\frac{\exp\left(w(\mathbf{u},\mathbf{v})\right)}{\sum_{\mathbf{v}^{\prime}\in\mathcal{N}(\mathbf{u})}\exp\left(w(\mathbf{u},\mathbf{v}^{\prime})\right)}$ (6)
-where $\mathcal{N}(\mathbf{u})$ denotes the set of nodes adjacent to $\mathbf{u}$ in the topic sub-graph.
-After obtaining the sampled topics, we next sample their corresponding key concepts.
-We first select an initial key concept by performing a one-step random walk on the topic-KC sub-graph, with transition probabilities computed using Equation (6) but applied specifically within this sub-graph.
-Subsequently, additional related key concepts are sampled through three to four steps of random walks on the KC sub-graph.
-This process results in a set of sampled topics and key concept $\mathbf{K}^{g}$.
-Since the sampled or extracted high-level concepts essentially serve as a summary, they may omit certain details originally present in the reference documents.
-Thus, we search for reference documents closely aligned with our sampled concept set $\mathbf{K}^{g}$.
-Specifically, we calculate the Jaccard similarity between the concept set $\mathbf{K}^{g}$ and the concept sets of each document within the filtered seed document set $\mathcal{D}$.
-Subsequently, we select the two documents with the highest similarity scores as the final reference documents.
-Analogous to the procedure in the Level-2 Generator, we instruct $\mathbf{M}^{Q}$ to randomly select and combine concepts from $\mathbf{K}^{g}$, and then utilize the selected concepts as well as the two chosen documents as references for question generation.
-The detailed prompt is shown in Figure 10 of Appendix.
-In the Level-3 Generator, the combination of concepts go beyond a single document.
-In addition to grounding on multiple documents, by employing random sampling on a concept graph constructed from the entire filtered document set, the diversity of the generated questions is significantly increased, thereby broadening the scope and richness of the resulting question set.
+where $\mathcal{N}(\mathbf{u})$ denotes the set of adjacent nodes to $\mathbf{u}$ in the topic sub-graph.
+Next, we sample key concepts corresponding to the selected topics.
+We begin with a one-step random walk on the topic-KC sub-graph, using transition probabilities from Equation (6).
+Additional related key concepts are then sampled through three to four steps of random walks within the KC sub-graph, resulting in a sampled set $\mathbf{K}^{g}$.
+Since $\mathbf{K}^{g}$ serves as a high-level summary, certain details from the original reference documents may be omitted.
+Thus, we search for reference documents most aligned with $\mathbf{K}^{g}$ as grounding.
+Specifically, we compute the Jaccard similarity between $\mathbf{K}^{g}$ and the concept sets of each document in $\mathcal{D}$, selecting top two documents with the highest similarity scores as references.
+Following the Level-2 procedure, we instruct $\mathbf{M}^{Q}$ to randomly select and combine concepts from $\mathbf{K}^{g}$, using the selected concepts along with the two reference documents for question generation.
+The detailed prompt is provided in Figure 9 of Appendix.
+Unlike Level-2, which is limited to a single document, Level-3 grounds on multiple documents and employs concept sampling from a global graph, significantly enhancing the diversity and breadth of generated questions.
 3.3 Answer Generation
-Using methods described in Section 4.2, we obtain a large set questions.
-Now we are ready to generation answers.
-We employ open-source LLMs such as Qwen2.5-Math-72B-Instruct as our answer generator $\mathbf{M}^{A}$.
-We have not yet incorporated an additional answer verification process, since preliminary experiments indicate that our synthetic data has already yielded satisfactory results.
-We do believe that dedicating effort to constructing and validating answers (e.g., majority voting and multi-agent debating) can further enhance the quality of our synthetic data at additional computational cost.
-We plan to leave the answer verification process in further work.
+Using the methods described in Section 3.2, we obtain a large set of questions and proceed with answer generation.
+We employ open-source LLMs, such as Qwen2.5-Math-72B-Instruct, as our answer generator $\mathbf{M}^{A}$.
+Currently, we do not incorporate an additional answer verification process, as preliminary experiments indicate that our synthetic data already achieves satisfactory quality.
+The detailed experimental results are shown in Section C.3.
+However, further improvements could be made by refining answer validation through techniques like majority voting (Wang et al., 2022; Jiao et al., 2024) or multi-agent debate (Du et al., 2024; Khan et al., 2024), albeit at additional inference computational cost.
+We leave this for future work.
 4 Experiments
 4.1 Experimental Settings
-Models.
-For $\mathbf{M}^{Q}$, we utilize Mistral-Large-Instruct-2407 [45], as we believe that larger models inherently possess the essential internal knowledge required for effective concept extraction and question synthesis.
-For the answer generator, $\mathbf{M}^{A}$, we employ Qwen2.5-Math-72B [52].
-For trained models, we adopt the Llama-3.2-1B, Llama-3.2-3B, and Llama-3.1-8B base models [31].
-Training Settings.
-We employ Supervised Fine-Tuning (SFT) to train the model.
-We adopt AdamW optimizer [33] and set learning rate to $1\times 10^{-5}$ with a batch size of $512$.
-We utilize the linear learning rate scheduler and train the model for 3 epochs.
-Our Datasets.
-[…]
-Statistics of Synthetic Questions and Answers.
-The median lengths of the questions generated by Level-1, Level-2, and Level-3 are $32$, $66$, and $80$, respectively, while the corresponding median lengths of the answers are $358$, $497$, and $545$.
-Evaluation.
-To evaluate the math reasoning ability of our model trained on synthetic data, we utilize the evaluation code from Qwen-Math [52] to conduct assessments on several mainstream benchmarks: MATH [15], GSM8K [10], OlympiadBench-Math (Olympiad) [14], CollegeMath (College) [43], Minerva Math (Minerva) [22], and Gaokao 2023 En (Gaokao) [28].
-Other Baselines.
-We compare our synthetic data with other high-quality synthetic datasets.
-MAmmoTH2 [55], NaturalReasoning [54], and JiuZhang3.0 [56] employ a direct extraction-based synthesis approach on pre-training documents, similar to our Level-1 method.
-OpenMathInstruct-2 [47] is a large-scale synthetic math dataset consisting of 14M samples, with GSM8K and MATH serving as the seed datasets.
-Additionally, we compare our dataset with NuminaMath [25], a high-quality dataset that includes both synthetically generated data and extracted question-answer pairs from exam papers and mathematics discussion forums.
-Furthermore, we evaluate our Level-2 and Level-3 approaches against existing methods designed to enhance question diversity through augmentation.
-Specifically, we include two popular augmentation approaches: the rephrase augmentation approach [53, 35] and the persona augmentation approach [12, 21], both of which generate new questions based on the questions extracted using Level-1.
-For rephrase augmentation, we follow the same procedure as MetaMath [53], providing two-shot examples in the prompt to rephrase questions.
-For persona augmentation, we utilize the 200K personas released by the authors as prompts for question augmentation.
-5.2 Main Results
-In this section, we present main evaluation results of SynthLLM.
-We test our models on various mathematical reasoning benchmarks, which encompass a range of difficulty levels.
-The results are shown in Table 2.
-The values in parentheses represent the sample size.
-For compared baseline training datasets, we directly evaluate model checkpoints they have released.
-We adopt the entire dataset synthesized with different methods from our framework to train models and denote it as SynthLLM-3B or -8B.
-The full dataset has a sample size of 7.4M, and we also create a subset with 3.2M samples.
-SynthLLM achieves superior performance compared to other synthetic datasets.
-SynthLLM consistently outperforms other synthetic datasets across all benchmarks, demonstrating the effectiveness of our framework.
-OpenMathInstruct-2, a high-quality math reasoning dataset generated using the Llama3.1-405B model based on MATH and GSM8K, comprises 14M samples.
-Despite its significantly larger size, the model trained on smaller dataset (SynthLLM-8B (3.2M)) achieves comparable performance on its in-distribution test sets (MATH and GSM8K) while substantially surpassing OpenMathInstruct-2 on Out-of-Distribution (OOD) test sets.
-This highlights the superior generalization capabilities of SynthLLM.
-Compared to the corresponding Llama Instruct models across different sizes, our models consistently achieve superior performance.
-Furthermore, our models (SynthLLM-8B) match or even surpass much larger models such as NuminaMath-CoT-72B and Llama-3.1-70B-Instruct on benchmarks, further reinforcing the effectiveness of our approach.
+Models and Our Datasets.
+For $\mathbf{M}^{Q}$, we use Mistral-Large-Instruct-2407 (team, 2024), as larger models inherently possess the internal knowledge necessary for effective concept extraction and question generation.
+For $\mathbf{M}^{A}$, we employ Qwen2.5-Math-72B (Yang et al., 2024).
+The trained models are Llama-3.2-1B, -3B, and Llama-3.1-8B base models (Llama Team, 2024).
+After filtering, we collected approximately $520\text{K}$ math documents.
+For Level-1 and Level-2, we generated 5 questions per reference document, yielding approximately $2.3\text{M}$ and $2.6\text{M}$ samples, respectively.
+For Level-3, each sampled $\mathbf{K}^{g}$ generated 3 questions, resulting in $2.6\text{M}$ samples.
+Further details of training and dataset are provided in Section B.
+Other Baselines and Evaluation Benchmarks.
+We compare our approach with MAmmoTH2 (Yue et al., 2024), JiuZhang3.0 (Zhou et al., 2024), NaturalReasoning (Yuan et al., 2025), and OpenMathInstruct-2 (Toshniwal et al., 2025).
+Additionally, we evaluate our Level-2 and Level-3 methods against existing augmentation techniques aimed at enhancing question diversity.
+Specifically, we consider two widely used approaches: rephrase augmentation (Yu et al., 2024) and persona augmentation (Ge et al., 2024; Lambert et al., 2024; Wang et al., 2025), both of which generate new questions based on those extracted in Level-1.
+Following Qwen-Math (Yang et al., 2024), we adopt evaluation benchmark: MATH (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), OlympiadBench-Math (Olympiad) (He et al., 2024), CollegeMath (College) (Tang et al., ), Minerva Math (Minerva) (Lewkowycz et al., 2022), and Gaokao 2023 En (Gaokao) (Liao et al., 2024).
+The more details are shown in Section B.
+4.2 Main Experimental Results
+We evaluate SynthLLM across multiple mathematical reasoning benchmarks of varying difficulty levels.
+Table 2 presents results, with values in parentheses indicating sample sizes.
+For baseline datasets, we directly evaluate their released models.
+We train models using both the full SynthLLM dataset (7.4M samples) and a subset obtained by uniformly sampling 3.2M examples, denoted as SynthLLM-1B, SynthLLM-3B, and SynthLLM-8B.
+SynthLLM Outperforms Other Synthetic Datasets.
+SynthLLM consistently surpasses other synthetic datasets across all benchmarks, demonstrating the effectiveness of SynthLLM.
+OpenMathInstruct-2 is a math dataset curated from MATH and GSM8K.
+Despite its larger size, the model trained on smaller dataset (SynthLLM-8B (3.2M)) achieves comparable performance on in-distribution test sets (MATH and GSM8K) while significantly outperforming it on OOD benchmarks, highlighting the superior generalization of SynthLLM.
+Additionally, our models consistently surpass Llama Instruct models of similar sizes and even match or outperform much larger models, such as NuminaMath-CoT-72B and Llama-3.1-70B-Instruct.
+Notably, we currently generate only five questions per document for Level-1 and Level-2 and three questions for Level-3.
+As shown in Figure 5, this is far from fully utilizing each document’s potential.
+Simply increasing the number of synthesized questions per document could further enhance benchmark performance.
 Model GSM8K MATH Minerva Olympiad College Gaokao Average
@@
 JiuZhang3.0-8B (6M) $88.7$ $51.2$ $21.7$ $18.8$ $37.5$ $43.4$ $43.6$
-NaturalReasoning-8B (2.8M)* - $55.6$
+NaturalReasoning-8B (2.8M)* - $55.6$ - - - - -
 OpenMathInstruct-2-8B (14M) $91.1$ $67.5$ $22.5$ $27.7$ $39.2$ $53.5$ $50.3$
@@
 NaturalReasoning-8B: Results are directly referenced from the original paper, as the full dataset and model have not been released.
-More importantly, for each reference document, we currently generate only five questions per document for Level-1 and Level-2, and three questions for Level-3, which is far from reaching the full potential of our method’s document utilization (as shown in Figure 6).
-We can further enhance performance on benchmarks simply by increasing the number of synthesized questions per document.
 4.3 Ablation Studies about Our SynthLLM Framework
-In this section, we present detailed ablation studies about our Level-2 and Level-3 methods, demonstrating their advantages in scalability and diversity.
+Comparison with Direct Extraction-Based Synthesis.
+We compare Level-1 (direct extraction-based synthesis) with Level-2 and Level-3 across models of different sizes (1B and 3B).
+As shown in Table 4 of Appendix, on the 3B model, Level-2 and Level-3 significantly outperform Level-1 and the instruct-tuned version across all benchmarks.
+On the 1B model, Level-3 achieves substantial gains over Level-1 on most benchmarks, except for College and Gaokao.
+While each document yields five sampled questions, the final dataset size for Level-1 (2.3M) is notably smaller than that of Level-2 (2.6M).
+This is due to the high redundancy in direct extraction, where many duplicate questions are generated and later removed.
+This highlights the scalability limitations of direct extraction-based approach.
+Diversity of Generated Questions.
+Compared to direct extraction-based synthesis, SynthLLM generates more diverse questions by decomposing and recombining knowledge concepts.
+To assess this, we compare Level-1 and Level-2 by measuring the similarity among generated questions within the same document.
+Specifically, we randomly select 2,000 documents and generate 50 questions per document for each method.
+Figure 4: Histogram of question similarity within the same document.
+The x-axis represents the cosine similarity between question pairs from the same document, while the y-axis denotes the frequency of each bin.
+We then use all-MiniLM-L6-v2 (Wang et al., 2020) to obtain question embeddings and compute cosine similarity between each question pair from same $d$.
+The histogram of question similarity is shown in Figure 4.
+The results clearly indicate that Level-2 generates more diverse questions, as evidenced by the lower similarity scores among questions derived from same $d$.
+Comparison with data augmentation methods on limited reference documents.
+We compare our Level-2 and Level-3 methods with augmentation techniques, including rephrase (Yu et al., 2024) and persona augmentation (Ge et al., 2024).
+We fix number of reference documents and evaluate different methods for scaling high-quality question generation from it.
+The effectiveness of each method will be measured by its ability to consistently improve model performance on downstream tasks as the dataset expands.
+Specifically, we randomly select 2,000 reference documents and generate 25, 50, to 150 questions per document using Level-2 and Level-3 methods, resulting in datasets ranging from 50K, 100K, to 300K questions.
+For rephrasing and persona augmentation, we need seed questions to conduct data augmentation.
+Therefore, we first adopt our Level-1 method to extract existing questions from the 2,000 fixed reference documents (yielding 6,500 questions) as the seed dataset.
+For above questions, we employ the same answer generator $\mathbf{M}^{A}$ (see Section 4.1).
+We train these datasets on the Llama-3.1-8B and present the results in Figure 5, where (a) shows MATH accuracy and (b) depicts average accuracy across benchmarks.
+Level-2 and Level-3 consistently outperform augmentation-based methods.
+Notably, as the number of generated questions per document increases to 150, our methods continue to yield performance gains, whereas rephrase and persona augmentation reach saturation.
+This highlights our approaches can efficiently utilize limited reference documents by decomposing and recombining knowledge concepts, enabling more scalable and high-quality question generation.
+In previous experiments, we generated only 5 questions per document for Level-2 and 3 questions per document for Level-3.
+These results suggest that our approach has not yet reached its full potential in document utilization and can be further expanded.
+Figure 5: (a): The performances of Level-2, Level-3, and other augmentation methods on MATH benchmark.
+(b): The average performances across various benchmarks.
+The x axis denotes the sample number.
+The y axis represents the accuracy (%).
+4.4 Preliminary Experiments in Coding Domain
+To explore the generalizability and effectiveness of our method, we conduct preliminary experiments in the coding domain.
+We adopt the same SYNTHLLM framework (described in Section 3) and follow the experimental setup outlined in Section 4.1.
+Specifically, we generate 2.2M samples using our Level-1 baseline method, and 2.4M and 2.5M samples for our Level-2 and Level-3 methods, respectively.
+Mistral-Large-Instruct-2407 is used as the question generator, while Qwen2.5-Coder-32B-Instruct serves as the answer generator.
+Model HumanEval MBPP
+Llama-3.1-8B-Instruct $72.6$ $71.2$
+Deepseek-Coder-v1.5-Instruct $75.6$ $73.4$
+SynthLLM Level-1 $73.2$ $67.3$
+SynthLLM Level-123 $78.7$ $72.5$
+Table 3: HumanEval and MBPP Performance (%)
+We then train the Llama-3.1-8B base model on these datasets.
+The results on HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) are summarized in Table 3.
+For comparison, we also include the performance of the instruction-tuned model (Llama-3.1-8B-Instruct) and a code-specific model (Deepseek-Coder-v1.5-Instruct (Guo et al., 2024)).
+SynthLLM Level-1 represents the model trained solely on Level-1 data, while the Level-123 model denotes the model trained on the entire synthetic dataset (combined data from 3 methods).
+Our newly proposed Level-2 and Level-3 methods yield significant improvements in coding performance over the Level-1 baseline.
+SynthLLM Level-123 outperforms the instruction-tuned model and achieves performance comparable to the code-specific model, demonstrating the generalizability and effectiveness of our approach in coding.
+5 Conclusions and Future Work
+We investigate scaling laws of synthetic data by introducing SynthLLM.
+Our findings show that synthetic data generated by SynthLLM follows a rectified scaling law, enabling accurate performance prediction when doubling the dataset size.
+This suggests that systematically expanding the synthetic dataset yields sustained and predictable performance gains.
+Comparisons with existing synthetic data generation and augmentation methods demonstrate that SynthLLM offers superior performance and scalability.
+Our framework could extend to domains such as physics, law, and finance.
+Future directions include refining SynthLLM to optimize pre-training data utilization and answer verification for improved accuracy.
+Appendix A Related Work
+Synthetic Data Generation.
+Synthetic data has emerged as a promising method to mitigate data scarcity and reduce the costs associated with data collection and annotation.
+In this work, we focus specifically on synthetic data generation for the post-training phase.
+Conventional methods typically leverage a limited number of high-quality human-annotated seed examples to generate diverse augmentations via large language models (LLMs).
+Common techniques include: 1) sampling seed instructions as few-shot prompts for generating new instructions (Wang et al., 2022; Honovich et al., 2022; Toshniwal et al., 2025; Li et al., 2024a); and 2) rephrasing seed samples to create variations (Yu et al., 2024; Maini et al., 2024; Ge et al., 2024; Lambert et al., 2024; Luo et al., 2023).
+However, reliance on scarce high-quality seeds limits scalability and diversity.
+Recent approaches by Xu et al.
+(2024) and Li et al.
+(2024b) propose generating synthetic data from scratch by exploiting either LLM uncertainty or predefined knowledge taxonomies.
+In contrast, pre-training data—vast and inherently diverse—remains underutilized for scalable post-training synthetic data generation.
+Prior methods have either directly extracted samples from web documents (similar to our Level-1 approach) (Yue et al., 2024; Zhou et al., 2024; Yuan et al., 2025) or used document backtranslation (Li et al., 2024).
+Our proposed SynthLLM method aligns with this research direction but provides a more efficient framework for leveraging web documents to scale the diversity of synthetic data.
+Scaling Law of LLMs.
+Scaling laws predict how model performance varies with model size and data volume, providing insights critical for optimizing computational resources during model training (Hestness et al., 2017; Rosenfeld et al., 2019; Kaplan et al., 2020; Hoffmann et al., 2022; Bahri et al., 2024).
+Recent advancements have introduced refined scaling laws, such as data-constrained scaling (Muennighoff et al., 2023), hyperparameter scaling Bi et al.
+(2024), and model distillation Busbridge et al.
+(2025).
+For fine-tuning LLMs, Hernandez et al.
+(2021) explored scaling behavior transitioning from unsupervised pre-training, emphasizing pre-training’s critical role.
+Lin et al.
+(2024) further introduced a rectified scaling law tailored specifically for fine-tuning tasks.
+Our work extends this rectified scaling law (Lin et al., 2024), demonstrating for the first time that scaling laws also apply robustly when fine-tuning LLMs using synthetic data.
+Appendix B Details of Training Model and Our Dataset
+Training Settings.
+We employ Supervised Fine-Tuning (SFT) to train the model.
+We adopt AdamW optimizer (Loshchilov & Hutter, 2019) and set learning rate to $1\times 10^{-5}$ with a batch size of $512$.
+We utilize the linear learning rate scheduler and train the model for 3 epochs.
+Our Dataset.
+In this work, we focus on the math domain.
+After filtering high-quality documents from web data, we collect approximately $520\text{K}$ math-related documents.
+For Level-1 and Level-2, we generate 5 questions per document, followed by deduplication and decontamination (Yue et al., 2024; Toshniwal et al., 2025) using evaluation benchmarks, yielding around $2.3\text{M}$ and $2.6\text{M}$ samples, respectively.
+For Level-3, the constructed concept graph contains around 32 00032\,000 topics and 200 000200\,000 knowledge concepts.
+We conduct random walks on the topic sub-graph for five epochs, covering each topic in each epoch.
+For each sampled $\mathbf{K}^{g}$, we generate 3 questions, producing about $2.6\text{M}$ samples.
+Due to the high computational cost of training on large-scale datasets, we limit the number of sampled questions per document in Level-2 and constrain the number of random walk epochs in Level-3 to a relatively small scale (e.g., five epochs).
+It is still far from fully leveraging the potential of each document (see Figure 5).
+For $\mathbf{M}^{Q}$, the output temperature is $0.75$.
+For $\mathbf{M}^{Q}$, the output temperature of $\mathbf{M}^{A}$ is 0.
+Statistics of Synthetic Questions and Answers.
+The median lengths of the questions generated by Level-1, Level-2, and Level-3 are $32$, $66$, and $80$, respectively, while the corresponding median lengths of the answers are $358$, $497$, and $545$.
+Other Baselines.
+We compare our synthetic data with other high-quality synthetic datasets.
+MAmmoTH2 (Yue et al., 2024), NaturalReasoning (Yuan et al., 2025), and JiuZhang3.0 (Zhou et al., 2024) employ a direct extraction-based synthesis approach on pre-training documents, similar to our Level-1 method.
+OpenMathInstruct-2 (Toshniwal et al., 2025) is a large-scale synthetic math dataset consisting of 14M samples, with GSM8K and MATH serving as the seed datasets.
+Additionally, we compare our dataset with NuminaMath (LI et al., 2024), a high-quality dataset that includes both synthetically generated data and extracted question-answer pairs from exam papers and mathematics discussion forums.
+Furthermore, we evaluate our Level-2 and Level-3 approaches against existing methods designed to enhance question diversity through augmentation.
+Specifically, we include two popular augmentation approaches: the rephrase augmentation approach (Yu et al., 2024; Maini et al., 2024) and the persona augmentation approach (Ge et al., 2024; Lambert et al., 2024), both of which generate new questions based on the questions extracted using Level-1.
+For rephrase augmentation, we follow the same procedure as MetaMath (Yu et al., 2024), providing two-shot examples in the prompt to rephrase questions.
+For persona augmentation, we utilize the 200K personas released by the authors as prompts for question augmentation.
+Evaluation.
+To evaluate the math reasoning ability of our model trained on synthetic data, we utilize the evaluation code from Qwen-Math (Yang et al., 2024) to conduct assessments on several mainstream benchmarks: MATH (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), OlympiadBench-Math (Olympiad) (He et al., 2024), CollegeMath (College) (Tang et al., ), Minerva Math (Minerva) (Lewkowycz et al., 2022), and Gaokao 2023 En (Gaokao) (Liao et al., 2024).
+The output temperature for all tested models is set to 0.
+Computational Cost.
+For generating questions, answers, and key concepts, we utilized a total of 32 Nvidia H100 GPUs, with each inference task running on 4 Nvidia H100 GPUs.
+The complete dataset generation process took approximately 3 days.
+For model training, we employed AMD MI300 GPUs; the largest training task utilized 32 AMD MI300 GPUs and required around 2 days to complete.
+Appendix C More Results
+In this section, we present additional experimental results about Section 4.
+C.1 Ablation Study about Our Three Distinct Methods
+We compare Level-1 (direct extraction-based synthesis) with Level-2 and Level-3 across models of different sizes (1B and 3B).
+As shown in Table 2 (Appendix), on the 3B model, Level-2 and Level-3 consistently outperform both Level-1 and the instruct-tuned version across all benchmarks.
+On the 1B model, Level-3 achieves substantial improvements over Level-1 on most benchmarks, except for College and Gaokao.
+Although each document contributes five sampled questions, the final dataset size for Level-1 (2.3M) is significantly smaller than that of Level-2 (2.6M).
+This discrepancy arises from the high redundancy in direct extraction, where many duplicate questions are generated and subsequently removed.
+These findings highlight the scalability limitations of direct extraction-based synthesis.
 Model GSM8K MATH Minerva Olympiad College Gaokao Average
@@
 Level-3-3B (2.6M) $77.0$ $49.1$ 18.0 17.2 37.4 $42.9$ 40.3
-Table 3: Performance of 1B and 3B models on various math benchmarks. The number in ()()( ) represents the sample size. The best result is in bolded. All metrics are measured in percentage (%).
-Comparison of Level-2 and Level-3 with directly extraction-based synthesis method.
-We compare the directly extraction-based synthesis method (Level-1) with our Level-2 and Level-3 methods across models of varying sizes.
-The results are shown in Table 3.
-We observe that, on the 3B model, Level-2 and Level-3 significantly outperform Level-1 and instruct version across benchmarks.
-On the 1B model, Level-3 surpasses Level-1 by a large margin on most benchmarks, with the exception of College and Gaokao.
-Although each document is sampled with 5 questions, the sample size for Level-1 (2.3M) is significantly lower than that of Level-2 (2.6M).
-Figure 5: Histogram of question similarity within the same document.
-The x-axis represents the cosine similarity between question pairs from the same document, while the y-axis denotes the frequency of each bin.
-This is because Level-1 is based on direct extraction, which leads to a substantial amount of duplicate questions being generated.
-After removing duplicates, the sample size decreases considerably.
-This also highlights the scalability limitations of direct extraction-based synthesis, as it is inherently constrained by the availability of high-quality reference documents containing questions.
-Diversity of Generated Questions.
-Compared to direct extraction-based synthesis, SynthLLM generates more diverse questions by decomposing and recombining knowledge concepts.
-To assess this, we compare Level-1 and Level-2 by measuring the similarity among generated questions within the same document.
-Specifically, we randomly select 2,000 documents and generate 50 questions per document for each method.
-We adopt all-MiniLM-L6-v2 [49] to obtain question embeddings and compute cosine similarity between each pair of questions within the same document.
-The histogram of question similarity is shown in Figure 5.
-We could clearly observe that questions generated by Level-2 exhibit lower similarity, indicating greater diversity among questions derived from the same document.
-Figure 6: (a): The performances of Level-2, Level-3, and other augmentation methods on MATH benchmark.
-(b): the average performances across various benchmarks.
-The x axis denotes the sample number.
-The y axis represents the accuracy.
-Comparison with other data augmentation methods on limited reference documents.
-We compare our Level-2 and Level-3 methods with mainstream augmentation techniques, including rephrase [53] and persona augmentation [12].
-We restrict the number of reference documents to a very small set and evaluate different methods for scaling high-quality question generation from a constrained document pool.
-The effectiveness of each method will be measured by its ability to consistently improve model performance on downstream tasks as the dataset expands.
-To achieve this goal, We randomly select 2,000 documents and generate 25, 50, 75, …, 150 questions per document—a significantly large number—using both Level-2 and Level-3 methods.
-This results in datasets of 50K, 100K, 150K, …, 300K questions for each method.
-For rephrase and persona augmentation methods, we first use Level-1 to extract existing questions from documents, yielding approximately 6,500 questions as seed dataset.
-We then expand this dataset to 300K questions using these two augmentation techniques.
-We train these datasets on Llama-3.1-8B model, with results presented in Figure 6.
-Figure 6 (a) illustrates MATH accuracy of different methods, while Figure 6 (b) shows average accuracy across benchmarks.
-Level-2 and Level-3 achieve significantly better performance on math benchmarks compared to augmentation methods.
-More importantly, as the number of generated questions per document increases to 150, our methods continue to show consistent improvements.
-In contrast, the performance of the rephrase and persona augmentation methods has already saturated as the dataset size expands.
-This demonstrates that compared with existing augmentation methods, our knowledge-guided methods more efficiently exploit limited reference documents by decomposing and recombining knowledge concepts, thereby enabling more scalable high-quality synthetic question generation.
-In previous experiments, we generated only 5 questions for Level-2 and 3 questions for Level-3.
-The results presented here indicate that our method has not yet reached its full potential in document utilization.
+Table 4: Performance of 1B and 3B models on various math benchmarks. The number in $()$ represents the sample size. The best result is in bolded. All metrics are measured in percentage (%).
+C.2 Ablation Study on Scaling Law Form
 Figure 6: The results of fitting scaling curves using different formulations on the Llama-3.2-3B model.
@@
 The green points represent the data sizes used to fit the scaling laws, while the red points are used to test the prediction performance of the fitted curves.
-5.4 Ablation Study on Scaling Law Form
 We compare the rectified form Eq.
@@
 This underscores the importance of introducing the pre-learned size $D_{l}$ for scaling law of fine-tuning LLMs.
-6 Conclusions and Future Work
-In this work, we investigate the scaling laws of synthetic data and propose SynthLLM, a scalable web-scale data synthesis framework that systematically transforms pre-training data into high-quality synthetic datasets.
-Our findings show that synthetic data generated through SynthLLM consistently adheres to a rectified scaling law across various model sizes.
-By leveraging the fitted scaling law, we accurately predict model performance when the synthetic dataset size is doubled.
-These results suggest that systematically expanding synthetic datasets can continue to drive sustained and predictable performance improvements.
-Furthermore, our comparisons with other synthetic data generation and augmentation methods demonstrate that SynthLLM achieves superior performance and better scalability, providing a more effective approach for further enhancing model performance.
-Our framework can be readily extended to other downstream domains, including code, physics and chemistry, and healthcare, expanding its applicability across diverse fields.
-Furthermore, SynthLLM is not limited to fine-tuning; we aim to explore its effectiveness in continued pre-training and even in the pre-training phase for synthetic data generation.
-Additionally, we will continue to refine and optimize our framework by developing more efficient strategies for leveraging pre-training data, with the goal of enhancing scaling efficiency of synthetic data.
-Appendix A Used Prompts
+C.3 Ablation Study on Answer Verification Process
+The decision not to conduct further answer verification was motivated by our prior preliminary experiments, which showed that removing noisy samples during SFT yielded only marginal gains in model performance.
+We describe these experiments below.
+Our preliminary experiments followed the same experimental setup as the ablation studies described in Section 4.3.
+We randomly selected 2,000 reference documents and generated 50 questions per document using our Level-2 method.
+For each question, we generated 8 candidate answers.
+Model MATH Average
+Llama-3.1-8B wo filtering $42.0$ $36.5$
+Llama-3.1-8B w filtering $42.2$ $36.6$
+Table 5: MATH and Average Performance (%)
+For each question, we used Llama-3.1-70B-Instruct as an answer evaluator to score its 8 candidate answers and discarded any judged "potentially incorrect".
+We then randomly picked one answer from the remaining candidates.
+Without filtering, one answer was randomly selected from all 8 candidates.
+In both cases, only one answer per question was retained for comparison.
+We then trained the Llama-3.1-8B base model on these generated datasets.
+We evaluate each trained model on the MATH benchmark and compute the average accuracy across all other benchmarks.
+As shown Table 5, filtering out noisy samples results in only marginal gains in model performance.
+Recent studies (Tang et al., ; Li et al., 2025; Gandhi et al., 2025) report consistent findings.
+As shown in Table 7 of (Tang et al., ), adding a validation step for synthetic data does not improve performance on mathematical reasoning tasks.
+Table 2 of (Li et al., 2025) and Figure 6 of (Gandhi et al., 2025) further show that models trained even on entirely incorrect answers achieve performance comparable to those trained on correct answers.
+These results, along with our experiments, suggest that for SFT, model performance remains relatively robust to a certain degree of noise in answer correctness.
+Appendix D Used Prompts
 In this section, we present prompts used in our SynthLLM framework.
@@
 topic 2 </topic> <key_concept> Key Concepts: 1.
-topic 1:       1.1.
-key concept       1.2.
-key concept        2.
+topic 1:    $1.$ key concept    1.2.
+key concept    … 2.
 topic 2:    2.1.
@@
 Figure 10: Prompt for extracting topics and key concepts from the document. The {{ text }} is the placeholders for inserting reference document.
-Appendix B Examples about Extracted Topics and Key Concepts from Reference Document
+Appendix E Examples about Extracted Topics and Key Concepts from Reference Document
 In this section, we demonstrate some examples about extracted topics and key concepts from reference document by using previous prompt shown in Figure 10.
 We adopt a relative large Language model, Mistral-Large-Instruct-2407.
-The examples are shown in Figure 12 and 13.
+The examples are shown in Figure 11 and 12.
 We can observe that Mistral-Large-Instruct-2407 is capable of extracting well-structured topics and their corresponding key concepts from the document.
@@
 Derivations and Proofs in Trigonometry </topic> <key_concept> Key Concepts: 1.
-Trigonometric Functions and Identities:       1.1.
-Sine, Cosine, and Tangent Functions       1.2.
+Trigonometric Functions and Identities:    $1.$ Sine, Cosine, and Tangent Functions    1.2.
 Trigonometric Identities (e.g., Pythagorean, Co-function, Sum and Difference)    1.3.
@@
 Geometry on a Sphere:    2.1.
-Latitude and Longitude       2.2.
-Circumference of a Circle on a Sphere       2.3.
+Latitude and Longitude    $2.$ Circumference of a Circle on a Sphere    2.3.
 Equation of a Circle on a Sphere    2.4.
@@
 Circular Motion and Angular Velocity    3.2.
-Calculating Angles in Right Triangles       3.3.
-Real-world Problems Involving Trigonometry (e.g., Crossing a Canyon)       3.4.
+Calculating Angles in Right Triangles    $3.$ Real-world Problems Involving Trigonometry (e.g., Crossing a Canyon)    3.4.
 Computing Sine and Cosine without a Calculator    3.5.
@@
 Complex Numbers and Trigonometric Functions    4.3.
-Roots of Unity and Their Relationship to Trigonometry       4.4.
-Inverse Hyperbolic Functions (e.g., Arcsinh)       4.5.
+Roots of Unity and Their Relationship to Trigonometry    $4.$ Inverse Hyperbolic Functions (e.g., Arcsinh)    4.5.
 Complex Numbers in Trigonometric Identities 5.
@@
 Derivation of Sum and Difference Formulas    5.4.
-Proof of the Haversine Formula       5.5.
-Derivation of Heron’s Formula </key_concept>
+Proof of the Haversine Formula    $5.$ Derivation of Heron’s Formula </key_concept>
 Figure 11: An example about High School’s Trigonometry.
@@
 Parametric Surfaces </topic> <key_concept> Key Concepts: 1.
-Surface Integrals of Vector Fields:       1.1.
-Definition of surface integral of a vector field       1.2.
+Surface Integrals of Vector Fields:    $1.$ Definition of surface integral of a vector field    1.2.
 Flux of a vector field across a surface    1.3.
@@
 Oriented Surfaces:    2.1.
-Definition of an oriented surface       2.2.
-Positive and negative orientations       2.3.
+Definition of an oriented surface    $2.$ Positive and negative orientations    2.3.
 Unit normal vectors and their role in orientation    2.4.
@@
 Definition and calculation of unit normal vectors    3.2.
-Gradient vector and its role in finding normal vectors       3.3.
-Normal vectors for surfaces given by z = f(x, y)       3.4.
+Gradient vector and its role in finding normal vectors    $3.$ Normal vectors for surfaces given by z = f(x, y)    3.4.
 Normal vectors for parametric surfaces    3.5.
@@
 Physical interpretation of flux in fluid dynamics    4.3.
-Calculation of flux using surface integrals       4.4.
-Flux across closed surfaces       4.5.
+Calculation of flux using surface integrals    $4.$ Flux across closed surfaces    4.5.
 Application of flux in Gauss’s Law 5.
@@
 Evaluation of surface integrals using parametric surfaces    5.4.
-Parameterization of common surfaces (e.g., spheres, cylinders)       5.5.
-Conversion between parametric and non-parametric forms </key_concept>
+Parameterization of common surfaces (e.g., spheres, cylinders)    $5.$ Conversion between parametric and non-parametric forms </key_concept>
 Figure 12: An example about the College’s Vector Calculus.